### v1

# To CoT or To Loop? A Formal Comparison Between Chain-of-Thought and Looped Transformers

Abstract
Chain-of-Thought (CoT) and Looped Transformers have been shown to empirically improve performance on reasoning tasks and to theoretically enhance expressivity by recursively increasing the number of computational steps. However, their comparative capabilities are still not well understood. In this paper, we provide a formal analysis of their respective strengths and limitations. We show that Looped Transformers can efficiently simulate parallel computations for deterministic tasks, which we formalize as evaluation over directed acyclic graphs. In contrast, CoT with stochastic decoding excels at approximate inference for compositional structures, namely self-reducible problems. These separations suggest the tasks for which depth-driven recursion is more suitable, thereby offering practical cues for choosing between reasoning paradigms.
1 Introduction
Transformer-based large language models (LLMs)(Vaswani et al., 2017) have achieved impressive performance across a wide range of tasks and have recently been extended to complex reasoning tasks. Rather than directly predicting final answers, prompting LLMs to generate intermediate reasoning steps, known as Chain-of-Thought (CoT) prompting(Wei et al., 2022), has been shown to significantly enhance reasoning capabilities. This has given rise to a broader paradigm in which inference-time compute is increasingly leveraged to support complex reasoning (Welleck et al., 2024).
This naturally raises the question: why is CoT prompting effective for complex reasoning tasks? Recent work has addressed this question by framing reasoning as a computational problem and analyzing its computational complexity. Feng et al. (2023) showed that certain classes of problems that cannot be solved with standard Transformers become solvable with CoT. Subsequent studies analyzed the complexity class of CoT with respect to the number of steps, compared to models such as Turing machines (Merrill & Sabharwal, 2024), Boolean circuits (Li et al., 2024), and probabilistic Turing machines (PTMs) (Nowak et al., 2024). In summary, these findings underscore that the strength of CoT lies in its ability to increase the effective number of computational steps.
Recently, Looped Transformer (Looped TF) was proposed as an architecture that recursively applies fixed-size Transformer layers by feeding their outputs back into their inputs (Dehghani et al., 2019). Similar to CoT, Looped TF increases the number of computational steps. However, they achieve this increase implicitly via architectural recursion rather than through explicit token-level decoding. The recursive structure has been shown to enhance expressivity: theoretical results establish Turing completeness(Giannou et al., 2023) and universal approximation (Xu & Sato, 2025), while empirical work demonstrates improved performance on reasoning tasks (Saunshi et al., 2025).
The existence of two distinct approaches raises a fundamental question:
What is the separation between Chain-of-Thought and Looped Transformer?
This question has motivated comparisons of their expressivity under the same number of steps or loops. Saunshi et al. (2025) showed an expressivity inclusion of CoT within TF, and Merrill & Sabharwal (2025) showed a separation favoring Looped TF in the logarithmic regime. While these results suggest a superiority of looped models over CoT, several important questions remain unanswered:
  • Does separation exist beyond the logarithmic regime? → Affirmative; see Section 4.
  • Are Looped Transformers universally more expressive than CoT? → Negative; see Section 5.
In this paper, we aim to clarify the distinct characteristics of Looped TF and CoT by analyzing two problem settings, thereby addressing these questions.
  • Looped TF for Parallel Computation: For deterministic settings, we analyze computational structures as directed acyclic graphs and reveal a fundamental difference: CoT is inherently sequential, whereas Looped TF enables parallel solutions. Furthermore, we establish a separation under the same polylogarithmic bound on the number of steps or loops.
  • Approximate Inference with CoT: We study the problem of approximately sampling from the uniform distribution over structured solutions, formalized via self-reducibility (Schnorr, 1976). The complexity of this task aligns with that of approximate counting, highlighting the advantage of probabilistic CoT over deterministic Looped TF. Moreover, we show that CoT is not only expressive, but that even a weak CoT model, when combined with self-consistency, can successfully solve the task.
2 Preliminaries and Notation
We consider models based on Transformer blocks that are applied recursively. Specifically, we analyze two architectures that differ in how intermediate computations are structured and reused across steps. We assume saturated attention, where attention weights are uniformly assigned to the tokens with the highest scores. For the feedforward computation, we consider either a standard fully connected layer or a mixture-of-experts (MoE) layer (Shazeer et al., 2017; Csordás et al., 2024). Formal definitions of the Transformer block are deferred to Appendix A.
2.1 Chain-of-Thought
A decoder-only Transformer with causal masking is used for next-token prediction. In CoT, the model is allowed to generate intermediate steps before producing the final answer. At each step, it decodes a single token conditioned on the entire current sequence, appends it to the input, and proceeds to the next step. The intermediate computation steps are explicitly represented as tokens.
Definition 2.1 (Informal: CoT).
Let ${\mathcal{V}}$ be a vocabulary, and let $f:{\mathcal{V}}^{*}\to{\mathcal{V}}$ denote an autoregressive decoder-only Transformer, where decoding is performed either deterministically or stochastically as a next-token prediction. Given an input sequence $x=(x_{1},\dots,x_{n})\in{\mathcal{V}}^{n}$, define the initial state as $f^{0}(x)\coloneq x.$ Then, for each step $k$, the CoT is defined recursively as
$f^{k+1}(x)\coloneq f^{k}(x)\cdot f(f^{k}(x)),$ (1)
where $\cdot$ denotes sequence concatenation. The final output after $T(n)$ steps is defined as the last $m$ tokens of $f^{T(n)}(x)$, corresponding to the answer.
2.2 Looped Transformer
Looped TF applies a non-causal Transformer block repeatedly to a fixed-length sequence of internal representations. The intermediate computations are maintained in the embedding space, rather than as explicit tokens, at each position and are iteratively refined across loop iterations.
Definition 2.2 (Informal: Looped TF).
Let $f:\mathbb{N}\times\mathbb{Q}^{d\times n}\to\mathbb{Q}^{d\times n}$ denote a standard Transformer layer without causal masking, where the first argument $k\in\mathbb{N}$ allows injecting loop-dependent positional encodings. Here, $\mathbb{Q}$ denotes the data type, and $d$ is the embedding dimension. Given an input sequence $x=(x_{1},\dots,x_{n})\in\mathcal{V}^{n}$, define the initial hidden state as $f^{0}(x)\coloneq(e(x_{1}),\dots,e(x_{n}))\in\mathbb{Q}^{d\times n},$ where $e:\mathcal{V}\to\mathbb{Q}^{d}$ is a token-wise embedding function. Then, for each loop $k$, the Looped TF is defined recursively as
$f^{k+1}(x)\coloneq f(k,f^{k}(x)).$ (2)
After $T(n)$ loop iterations, the final output is obtained by applying a linear projection and deterministic decoding (e.g., $\arg\max$) to each position in $f^{T(n)}(x)$, extracting the last $m\leq n$ tokens.
2.3 Notation
For any $n\in\mathbb{N}^{+}$, we write $[n]\coloneq\{1,2,\ldots,n\}$. For functions $f,g:\mathbb{N}\to\mathbb{N}$, we write $f(n)\in O(g(n))$ iff there exist constants $c>0$ and $n_{0}\in\mathbb{N}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$. We denote by $\mathsf{poly}(n)$ the set of functions that grow at most polynomially, i.e., $\mathsf{poly}(n)\coloneq\left\{f:\mathbb{N}\to\mathbb{N}\;\middle|\;\exists k\in\mathbb{N},\;\exists c>0,\;\forall n\in\mathbb{N},\;f(n)\leq c\cdot n^{k}\right\}.$ Given two complexity classes ${\mathcal{C}}_{1}$ and ${\mathcal{C}}_{2}$, we write ${\mathcal{C}}_{1}\setminus{\mathcal{C}}_{2}$ to denote the set of decision problems (referred to as languages in computational complexity theory) that belong to ${\mathcal{C}}_{1}$ but not to ${\mathcal{C}}_{2}$.
3 General Problem Setting
This section begins by reviewing the problem settings considered in prior work, as summarized in Table 1, and identifying the remaining open questions. We then formalize the two problem settings addressed in this study, specifying their underlying assumptions and objectives.
3.1 Related Work
Table 1: Summary of theoretical analyses on expressivity of CoT and Looped TF.
Paper Model Type Problem Setting Class
Prystawski et al. (2023) CoT Pro. Bayesian network -
Feng et al. (2023) CoT Det. Mathematics & Decision-making
Merrill & Sabharwal (2024) CoT Det. Automata & Turing machine
Li et al. (2024) CoT Det. Boolean circuit
Nowak et al. (2024) CoT Pro. Language modeling (PTM)
Saunshi et al. (2025) Looped Det. Non-looped, CoT, and Automata -
Merrill & Sabharwal (2025) Looped Det. Automata & Graph connectivity
Ours Looped Det. Directed acyclic graph
&CoT Pro. FPRAS & FPAUS (self-reducibility)
Expressivity of CoT
Prior analyses of CoT can be broadly categorized into deterministic and probabilistic settings, each corresponding to a distinct problem formulation and focusing on different aspects of expressivity. In the deterministic setting, Feng et al. (2023) showed that CoT enables solving certain function tasks, such as mathematical and dynamic programming problems. Subsequent works clarified the computational classes corresponding to different numbers of reasoning steps by drawing comparisons to formal models of computation: Merrill & Sabharwal (2024) analyzed CoT through its correspondence with automata and Turing machines, while Li et al. (2024) established a connection to Boolean circuits, illustrating how CoT’s expressivity scales with the number of reasoning steps. In contrast, the probabilistic setting focuses on the model’s ability to represent and reason over distributions. Prystawski et al. (2023) studied CoT with sampling-based decoding for probabilistic inference, and Nowak et al. (2024) further investigated its capacity to model distributions over strings in language modeling, establishing a connection to PTMs.
Expressivity of Looped TF
Looped TF inherently operates deterministically and has therefore primarily been studied on deterministic tasks, with their expressivity analyzed in comparison to deterministic CoT under an equal number of steps or loops. Saunshi et al. (2025) showed that a Looped TF with $T$ iterations can simulate $T$ steps of a non-looped or CoT, thereby demonstrating an inclusion of CoT within the expressivity of Looped TF. Furthermore, Merrill & Sabharwal (2025) established that Looped TF with only $\log n$ iterations can solve problems such as graph connectivity, which is $\mathsf{NL}$-complete, whereas nonuniform CoT with $\log n$ steps remain within $\mathsf{TC}^{0}$, thereby proving a strict separation in complexity class at the logarithmic scale.
3.2 Open Questions and Motivations
We now outline open questions in the existing literature, separately for deterministic and probabilistic settings, each of which motivates the problem formulations studied in this work.
  • Deterministic Setting: Although Looped TF has been studied on specific tasks, its overall computational characteristics remain unclear. In contrast to CoT, which has been characterized as sequential nature (Feng et al., 2023; Merrill & Sabharwal, 2024; Li et al., 2024), the computational properties of Looped TF remain unexplored, in particular, their relationship to parallel computation (Sanford et al., 2024). Furthermore, while the expressiveness of CoT has been analyzed across various step regimes, that of Looped TF beyond the logarithmic regime, as well as any formal separation from CoT, has yet to be established.
  • Probabilistic Setting: While probabilistic CoT has been investigated primarily for distribution modeling, its applicability to solving deterministic problems remains insufficiently understood, limiting direct comparisons with Looped TF. Moreover, there remains a gap between the theoretical capabilities of CoT as PTMs (Nowak et al., 2024) and the empirical behavior of practical LLMs: such models are typically trained to minimize loss up to a bounded error, and inference-time techniques are necessary to exploit their probabilistic behavior for improved performance.
3.3 Our Problem Settings
We consider two fundamental classes of computational problems: the deterministic problem and the probabilistic problem, elaborated in Section 4 and Section 5, respectively. This section provides high-level definitions that specify the input–output structure in each setting; precise formulations are deferred to the corresponding sections. These problem classes differ in several fundamental aspects, such as output uniqueness, the presence of a target distribution, and the inclusion of accuracy parameters.
Definition 3.1 (Deterministic Problem).
Let $f:\mathcal{X}\to\mathcal{Y}$ be a target function. A model $M$ solves the deterministic problem for $f$ if, for every input $x\in\mathcal{X}$, it holds that $M(x)=f(x)$.
Definition 3.2 (Probabilistic Problem).
Let $R\subseteq\mathcal{X}\times\mathcal{Y}$ be a relation such that each input $x\in\mathcal{X}$ may have multiple valid outputs defined by $R(x)\coloneq\{y\in\mathcal{Y}\mid(x,y)\in R\}.$ A target distribution $p(\cdot\mid x)$ is specified over the set $R(x)$. The probabilistic problem is defined with respect to this distribution and is evaluated against explicitly specified accuracy or failure parameters provided as part of the input.
4 Looped Transformer Enables Efficient Parallel Solutions
This section analyzes the fundamental computational differences between CoT and Looped TF in the deterministic setting. We formalize deterministic computations using computation graphs, as introduced in Section 4.1. In Section 4.2, we show how CoT and Looped TF differ in processing the same computation graphs, highlighting their respective strengths. Finally, Section 4.3 presents a separation in their computational complexity in the polylogarithmic regime.
4.1 Problem Setting
Figure 1: Comparison of simulating strategies for computing a DAG. (a) A computation graph. (b) A Looped TF simulates the graph in a layer-wise and parallel fashion, requiring a number of loops proportional to the depth of the graph. (c) A CoT simulates the computation sequentially in topological order, requiring a number of steps proportional to the size of the graph.
To analyze the behavior of CoT and looped models, we formalize deterministic computations as directed acyclic graphs (DAGs), where each node corresponds to a local function and each edge indicates a data dependency. This abstraction is natural and general, as straight-line programs can be expressed as DAGs (Aho & Ullman, 1972). An illustrative example is shown in Fig. 1(a).
Definition 4.1 (Computation Graph).
Let $\Sigma$ be a finite alphabet. A computation graph is a directed acyclic graph $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$ that defines a function $G_{n}:\Sigma^{n}\to\Sigma^{m}$, where $\mathcal{F}_{n}=\{f_{n}:\Sigma^{\ell_{n}}\to\Sigma\mid\ell_{n}\in\mathbb{N}\}$ is a finite set of local functions associated with each node. The graph $G_{n}$ consists of (i) input nodes with in-degree 00, (ii) function nodes labeled by $f\in\mathcal{F}_{n}$ whose in-degree matches the arity of the corresponding function, and (iii) $m$ output nodes $O\subseteq V_{n}$ with out-degree 00. The overall function is obtained by evaluating the graph in topological order. The size of the graph is $|V_{n}|$, and its depth is the length of the longest path from an input node to an output node. We assume that the size of the computation graph $|V_{n}|$ is bounded by a polynomial in the input length $n$.
Remark. Prior work has analyzed CoT from DAGs, but in specific instances: Feng et al. (2023) showed that dynamic programming tasks can be simulated via DAGs, and Li et al. (2024) related CoT to Boolean circuits, a subclass of DAGs. We extend this view by modeling general computations as DAGs, yielding a unified framework for analyzing both CoT and Looped TF and clarifying how they simulate such computations in fundamentally different ways.
4.2 Simulating Computation DAGs: CoT Needs Size-scaled Steps and Looped Transformers Need Depth-scaled Loops
To characterize the expressiveness of CoT and Looped TF, we analyze how each simulates computation graphs. We begin by formalizing the assumptions under which computations can be implemented by Transformer-based models (see Section B.1 for details).
Assumption 4.2 (Polynomially-Efficient Approximation (cf. Feng et al., 2023)).
Each function in the computation graph can be approximated by a feedforward neural network of polynomial size.
Assumption 4.3 (Bounded Fan-in).
The graph has a bounded fan-in.
We now formalize the expressivity of CoT in simulating DAGs under these assumptions:
Theorem 4.4 (CoT for DAGs).
Let $\{G_{n}\}_{n\in\mathbb{N}}$ be a family of computation graphs, each computing a function $G_{n}:\Sigma^{n}\to\Sigma^{m(n)}$, and satisfying Assumptions 4.2 and 4.3. Then, for each $n\in\mathbb{N}$, there exists a log-precision CoT with parameter size bounded by $\mathsf{poly}(n)$, such that for every input $x\in\Sigma^{n}$, the model outputs $G_{n}(x)$ in a number of steps proportional to the “size” of the graph $G_{n}$.
We then analyze Looped TF, which requires an additional assumption. Unlike CoT, Looped TF does not extend the input sequence; consequently, the amount of information that can be retained at each loop iteration is limited.
Assumption 4.5 (Linear Parallel Space Complexity).
The amount of memory that must be maintained concurrently to evaluate the computation graph in parallel is $O(n)$.
Assumption 4.6 (Output Bound).
The computation graph contains at most $n$ output nodes.
Under these constraints, we formalize the expressivity of looped models in simulating DAGs.
Theorem 4.7 (Looped TF for DAGs).
Let $\{G_{n}\}_{n\in\mathbb{N}}$ be a family of computation graphs, each computing a function $G_{n}:\Sigma^{n}\to\Sigma^{m(n)}$, and satisfying Assumptions 4.24.34.5, and 4.6. Then, for each $n\in\mathbb{N}$, there exists a log-precision Looped TF with parameter size bounded by $\mathsf{poly}(n)$, such that for every input $x\in\Sigma^{n}$, it computes $G_{n}(x)$ using a number of loops proportional to the “depth” of the graph $G_{n}$.
Remark. Proofs are given in Appendix B, with illustrations in Fig. 1. Informally, Theorem 4.4 shows that CoT simulates computations by decoding nodes in topological order, whereas Theorem 4.7 shows that Looped TF simulates computations in a layer-wise and parallel fashion. This contrast reflects complementary strengths: CoT employs intermediate steps as scratchpad memory, supporting general computation without structural constraints. In contrast, Looped TF exploits structural parallelism to achieve greater efficiency, but requires additional assumptions.
4.3 Separation between CoT and Looped Transformers in Polylogarithmic Time
To compare the computational complexity of CoT and Looped TF, we analyze them with respect to Boolean circuits, a model of computation represented by DAGs. In particular, we restrict our attention to languages. We begin by defining the complexity classes of CoT, as defined in Li et al. (2024), and Looped TF.
Definition 4.8 ($\mathsf{CoT}$ and $\mathsf{LOOP}$).
Let $\mathsf{CoT}[T(n),d(n),p(n)]$ (resp. $\mathsf{LOOP}[T(n),d(n),p(n)]$) denote the set of languages :${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a CoT (resp. Looped TF) $M_{n}$ for each input size $n$, with embedding size $O(d(n))$ and $O(p(n))$ bits of precision, such that for all $x\in\{0,1\}^{n}$, the final output token, after $O(T(n))$ decode steps (resp. loops), is ${\mathcal{L}}(x)$, i.e., $M_{n}(x)={\mathcal{L}}(x),\text{ for all }x\in\{0,1\}^{n}.$
To characterize the computational power of these models, we refer to standard nonuniform circuit complexity classes. In particular, we focus on the classes $\mathsf{NC}^{k}$ and $\mathsf{TC}^{k}$, sing Boolean and threshold circuits, respectively, which capture efficient parallel computation (Arora & Barak, 2009).
Definition 4.9 ($\mathsf{TC}^{k}$ and $\mathsf{NC}^{k}$).
For each integer $k\geq 0$, the nonuniform circuit classes are defined as:
  • $\mathsf{NC}^{k}$ is the class of languages decidable by uniform Boolean circuits of polynomial size and depth $O(\log^{k}n)$, with bounded fan-in AND, OR, and NOT gates.
  • $\mathsf{TC}^{k}$ is the class of languages decidable by uniform threshold circuits of polynomial size and depth $O(\log^{k}n)$, with unbounded fan-in majority (threshold) gates.
Since Boolean circuits allow polynomial-size, we assume a polynomial embedding size. Prior work has shown that CoT is equivalent to Boolean circuits composed of $\mathsf{TC}^{0}$ gates, with circuit “size” proportional to the number of CoT steps. This class $\mathsf{SIZE}^{\mathsf{TC}^{0}}$ is defined in Definition B.6.
Theorem 4.10 (Li et al., 2024).
$\forall T(n)\in\mathsf{poly}(n),\mathsf{CoT}[T(n),\mathsf{poly}(n),\log{n}]=\mathsf{SIZE}^{\mathsf{TC}^{0}}[T(n)+1]$
In contrast, we show that Looped TF corresponds to Boolean circuits whose “depth” is proportional to the number of loop iterations.
Theorem 4.11 (Looped and Circuits).
Let $T(n)\in\mathrm{poly}(n)$. Then, a decision problem ${\mathcal{L}}$ belongs to the class $\mathsf{LOOP}[T(n),\mathrm{poly}(n),\log n]$ if and only if ${\mathcal{L}}$ is decidable by nonuniform families of threshold circuits of polynomial size and depth $T(n)$. Similarly, ${\mathcal{L}}\in\mathsf{LOOP}[T(n),\mathrm{poly}(n),1]$ if ${\mathcal{L}}$ is decidable by nonuniform families of Boolean circuits of polynomial size and depth $T(n)$.
These results formally separate their expressive power under standard complexity assumptions.
Theorem 4.12.
$\mathsf{TC}^{k-1}\subsetneq\mathsf{NC}^{k}\Rightarrow\mathsf{CoT}[\log^{k}{n},\mathsf{poly}(n),\log{n}]\subsetneq\mathsf{LOOP}[\log^{k}{n},\mathsf{poly}(n),1].$
Theorem 4.13.
$\mathsf{TC}^{k-1}\subsetneq\mathsf{TC}^{k}\Rightarrow\mathsf{CoT}[\log^{k}{n},\mathsf{poly}(n),\log{n}]\subsetneq\mathsf{LOOP}[\log^{k}{n},\mathsf{poly}(n),\log{n}].$
Proof.
It suffices to show that $\mathsf{CoT}[\log^{k}n,\mathsf{poly}(n),\log n]\subseteq\mathsf{TC}^{k-1}$ for any $k$. For any input, a CoT with $\log n$ steps yields only $2^{\log n}=\mathsf{poly}(n)$ possible execution traces, each of which can be simulated in $\mathsf{TC}^{0}$. Thus, we divide the total $\log^{k}n$ steps into $\log^{k-1}n$ blocks, each consisting of $\log n$ steps. We then simulate each block in parallel using $\mathsf{TC}^{0}$, and iterate this process over $\log^{k-1}n$ layers. This layered simulation can be performed within $\mathsf{TC}^{k-1}$, completing the proof. ∎
Remark. These results show that, under the same polylogarithmic number of steps or loops, Looped TF can simulate parallel computations more efficiently than CoT. This highlights the inherent parallelism of looped models, in contrast to the sequential nature of CoT.
Instance Examples
There are several concrete problems in (nonuniform) $\mathsf{NC}^{2}$: solving linear equations, testing for perfect matching in bipartite graphs, and membership testing for fixed context-free grammars, denoted by LIN, MATCH, and FCFL, respectively (see Section B.6 for details). Then we have:
Proposition 4.14.
$\textsc{LIN},\textsc{MATCH},\textsc{FCFL}\in\mathsf{LOOP}[\log^{2}{n},\mathsf{poly}(n),1]\setminus\mathsf{CoT}[\log^{2}{n},\mathsf{poly}(n),\log{n}].$
5 Chain-of-Thought Enables Probabilistic Reasoning
In this section, we investigate the expressiveness of CoT with stochastic decoding for probabilistic tasks. Section 5.1 introduces the task formulation and defines the model. Section 5.2 analyzes the expressive power of CoT in an idealized setting and establishes a formal separation from Looped TF. Finally, Section 5.3 shows that even weak CoT, when combined with self-consistency, exhibits sufficient expressive power.
5.1 Problem Setting
Structured Reasoning as Self-Reducibility
Complex reasoning tasks, such as solving puzzles, proving theorems, or planning actions, often require producing a globally consistent output through a sequence of interdependent local decisions, each addressing a subproblem. To formalize this compositional structure, we adopt the notion of self-reducibility (Schnorr, 1976).
Definition 5.1 (Informal: Self-reducibility (Schnorr, 1976)).
A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is self-reducible if any solution $y=y_{1}\dots y_{n}$ to input $x$ can be constructed incrementally, such that a prefix $y_{1:\sigma(x)}$ determines a subproblem $\psi(x,y_{1:\sigma(x)})$ whose solution completes $y$. See Definition C.4 for a formal definition.
For example, consider the Boolean satisfiability problem (SAT). A satisfying assignment can be constructed incrementally, where each partial assignment defines a simpler subproblem over the remaining variables. This process illustrates the canonical form of self-reducibility.
Task Formulation
We formalize the computational task of approximately sampling a structured solution.
Definition 5.2 ($\varepsilon$-Approximate Sampling (FPAUS)).
Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation, and let $p(\cdot\mid x)$ denote the uniform distribution over the solution set $R(x)\coloneq\{y\mid(x,y)\in R\}$. The task of $\varepsilon$-approximate sampling, also known as almost uniform generation, is to, for any given error parameter $\varepsilon>0$, output a distribution $q(\cdot\mid x)$ such that for all $x\in\Sigma^{*}$,
$\|q(\cdot\mid x)-p(\cdot\mid x)\|_{\mathrm{TV}}\leq\varepsilon,$ (3)
where $\|\cdot\|_{\mathrm{TV}}$ denotes the total variation distance.
Probabilistic CoT Model
We consider a CoT equipped with sampling-based decoding, where stochastic branching at intermediate steps allows for diverse reasoning paths and autoregressively defines conditional distributions over output tokens. This formulation aligns with the behavior of LLMs, which explore a space of possible reasoning trajectories and produce varied outputs across different sampling runs.
Definition 5.3 (Probabilistic CoT).
Let $V$ be a finite vocabulary and $V^{*}$ its Kleene closure. A probabilistic CoT stochastically generates, given an input $x\in V^{*}$, a sequence of output blocks of the form
$\langle r_{1}\rangle\langle e\rangle\langle y_{1}\rangle\langle e^{\prime}\rangle\;\langle r_{2}\rangle\langle e\rangle\langle y_{2}\rangle\langle e^{\prime}\rangle\;\cdots\;\langle r_{m}\rangle\langle e\rangle\langle y_{m}\rangle\langle e^{\prime}\rangle,$
where each $r_{i}\in V^{*}$ is a reasoning path, $e,e^{\prime}\in V$ are special delimiter tokens, yielding the final output string $y_{1}\dots y_{m}$. The generation proceeds autoregressively via next-token prediction: for each $i\in[m]$, the model generates a reasoning step $r_{i}$ followed by an output token $y_{i}$, conditioned on the input $x$, previous outputs $y_{<i}$, and prior reasoning steps $r_{<i}$. We denote by $\pi$ the next-token conditional distribution defined by the model.
Approximate Counting
Approximate sampling is closely related to approximate counting (Jerrum et al., 1986). For a relation $R$, define $f(x)\coloneq\lvert R(x)\rvert$, the number of solutions $y$ such that $(x,y)\in R$. A Fully Polynomial-Time Approximation Scheme (FPTAS) is a deterministic algorithm that, for any input $x$ and $\varepsilon>0$, outputs $\hat{f}(x)$ such that $(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$ in time polynomial in $|x|$ and $1/\varepsilon$. A Fully Polynomial Randomized Approximation Scheme (FPRAS) is a randomized version that additionally takes $\delta>0$, and satisfies the same guarantee with probability at least $1-\delta$. See Appendix C for details.
5.2 CoT is More Expressive than Looped TF in Approximate Inference
Whereas the previous section showed the superiority of Looped TF over deterministic CoT, we now present a separation in the opposite direction. Since deterministic models cannot perform sampling, we consider approximate counting; for self-reducible relations, it is equivalent to approximate sampling (Jerrum et al., 1986), and thus serves as a proxy for evaluating probabilistic inference capabilities.
Theorem 5.4 (Jerrum et al., 1986).
For self-reducible relations, approximate counting and approximate sampling are polynomial-time equivalent.
In computational theory, approximate sampling or counting algorithms are formalized as PTMs that run in polynomial time (Jerrum et al., 1986). Prior work has shown that probabilistic CoT can simulate PTM.
Lemma 5.5 (Informal: CoT as PTM (Nowak et al., 2024)).
Any polynomial-time probabilistic Turing machine can be simulated by a probabilistic CoT, such that the model reproduces the same output distribution and halts within a polynomial number of decoding steps.
Let $\mathsf{FPRAS}$ and $\mathsf{FPTAS}$ denote the function classes admitting an FPRAS and an FPTAS, respectively. Some problems admit an FPRAS, but no FPTAS is currently known for them (Jerrum et al., 2004). Intuitively, allowing failure (but can be small for arbitrarily high confidence) makes such problems tractable via randomized methods. Assuming this gap, we identify the first task for which CoT is more powerful than Looped Transformer:
Theorem 5.6.
Assuming that $\mathsf{FPTAS}\neq\mathsf{FPRAS}$, there exists a relation $R$ such that a probabilistic CoT can approximate the count $|R(x)|$ within any desired relative error $\varepsilon>0$ with high probability in polynomial time, whereas no (deterministic) Looped TF can achieve the same approximation within a polynomial number of loop iterations.
5.3 Expressiveness of Weak CoT
While Lemma 5.5 analyzes an idealized setting where a CoT, viewed as a PTM, is assumed to receive explicit error parameters, such assumptions diverge from the behavior of practical LLMs. In practice, models are trained within bounded error, but do not offer explicit control over the accuracy of their outputs. In this section, we examine the expressiveness of weak CoT models, incorporating inference-time strategies such as majority voting. The central question is: Can weak CoT be approximate samplers?
5.3.1 Formal Setting for Weak CoT
We consider models trained to minimize the next-token prediction loss as follows.
Definition 5.7 (Token-Wise Cross-Entropy Loss under Teacher Forcing).
For each input $x\in\Sigma^{*}$, let $p(y_{i}\mid x,y_{<i})$ denote the target conditional distribution over the next token $y_{i}\in\Sigma$ at position $i$, and let $\pi(y_{i}\mid x,y_{<i},r_{i})$ denote the model’s prediction conditioned on a reasoning trajectory $r_{i}$. Then, the token-wise cross-entropy loss under teacher forcing is defined as
$\mathcal{L}_{\mathrm{CE}}(x,y_{<i},r_{i})\coloneqq\mathbb{E}_{y_{i}\sim p(\cdot\mid x,y_{<i})}\left[-\log\pi(y_{i}\mid x,y_{<i},r_{i})\right],$ (4)
where $r_{i}$ denotes a reasoning trajectory sampled from $\pi(r_{i}\mid x,y_{<i})$.
We assume that each reasoning path $r_{i}$ depends only on the preceding output tokens $y_{<i}$, and is conditionally independent of the previous reasoning steps $r_{<i}$.
Assumption 5.8 (Conditional Independence of Reasoning Paths).
For each position $i\in[m]$, the reasoning path $r_{i}$ is conditionally independent of prior reasoning steps given the input and preceding outputs: $\pi(r_{i}\mid x,r_{<i},y_{<i})=\pi(r_{i}\mid x,y_{<i}).$
We further assume access to repeated samples of reasoning paths:
Assumption 5.9 (Sampling Independence).
We assume access to an unbounded number of independent samples of reasoning paths, i.e., for each $i\in[m]$, one can sample $r_{i}^{(1)},r_{i}^{(2)},\dots\sim\pi(\cdot\mid x,y_{<i})$ independently.
Then, we define the convergence of pertaining.
Definition 5.10 ($(\alpha,\gamma)$-Weak Token-Wise Learning).
Let $\alpha:\mathbb{N}\to\mathbb{R}_{>0}$ be any positive function, and let $\gamma\in(0,\tfrac{1}{2})$. We assume that for each token position $i\in[m]$, and for any input $x\in\Sigma^{*}$ and prefix $y_{<i}$, the model satisfies the following token-wise conditional cross-entropy loss bound under teacher forcing:
$\Pr_{r_{i}\sim\pi(\cdot\mid x,y_{<i})}\left[\mathcal{L}_{\mathrm{CE}}(x,y_{<i},r_{i})\leq\alpha(|x|)\right]\geq\frac{1}{2}+\gamma.$ (5)
Remark. This definition formalizes a weak yet meaningful correctness criterion for token-wise predictions in reasoning models. The task is assumed to be sufficiently difficult that directly predicting $y_{i}$ from $x$ and $y_{<i}$ alone does not yield low cross-entropy. Introducing an intermediate reasoning step $r_{i}$ is thus necessary to compensate for this gap. It guarantees the model achieves sufficiently low cross-entropy loss with probability at least $\tfrac{1}{2}+\gamma$. This raises the following questions:
  • Can a weak CoT support inference with arbitrarily high confidence?
  • Can a weak CoT be transformed into an approximate sampler for any desired accuracy?
These questions are addressed in the subsequent sections, with formal proofs deferred to Appendix D and key ideas shown in Fig. 2.
Figure 2: Illustration of next-token prediction in CoT as approximate relative counting over a self-reducible relation. Given a SAT instance $x$, self-reducibility enables reduction to a subproblem $x^{\prime}$ based on the autoregressive prefix (e.g., $y_{1}=1$). The probability $\pi(y_{2}=1)$ is approximated via relative counting, with self-consistency aggregating predictions from multiple reasoning paths.
5.3.2 Confidence Amplification via Self-Consistency
We show that self-consistency(Wang et al., 2023) can enhance the confidence. While prior work(Wu et al., 2025) focused on classification tasks, we consider settings in which the output is a probability distribution.
Lemma 5.11 (Coordinate-wise Self-Consistency).
Suppose that a CoT model $\pi$ achieves $(\alpha,\gamma)$-weak token-wise learning with respect to the target distribution $p(y_{i}\mid x,y_{<i})$, for some function $\alpha:\mathbb{N}\to\mathbb{R}_{>0}$ and constant $\gamma\in(0,\tfrac{1}{2})$. Let $\delta\in(0,1)$. Fix any token position $i\in[m]$. Let $r_{i}^{(1)},\dots,r_{i}^{(T)}\sim\pi(\cdot\mid x,y_{<i})$ be $T$ independent reasoning paths, and let $\pi^{(t)}=\pi(Y_{i}\mid x,y_{<i},r_{i}^{(t)})\in\Delta^{|\mathcal{V}|}$ denote the predicted distribution at position $i$ for each path. Define the consensus prediction $\pi^{\star}\in\mathbb{R}^{|\mathcal{V}|}$ by taking the coordinate-wise median:
$\pi^{\star}(y)=\operatorname{median}\left(\pi^{(1)}(y),\dots,\pi^{(T)}(y)\right)\quad\text{for each }y\in\mathcal{V}.$ (6)
Then, for $T\geq\frac{2}{\gamma^{2}}\log\left(\frac{|\mathcal{V}|}{\delta}\right)$, with probability at least $1-\delta$, the self-consistent prediction satisfies
$\left|p(y_{i}\mid x,y_{<i})-\pi^{\star}(y_{i})\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\quad\text{for all }y_{i}\in\mathcal{V}.$ (7)
That is, self-consistency improves the reliability of predictions by aggregating multiple reasoning paths.
5.3.3 Approximate Sampling via Self-Reducibility
We then show that the error of weak CoT can be corrected, yielding an approximate sampler. We begin by noting that next-token predictions become well-defined tasks when the underlying relation is self-reducible, thus justifying the realizability assumption underlying weak token-wise learning, as shown in Fig. 2.
Proposition 5.12 (Reducing Next-Token Prediction to Relative Counting via Self-Reducibility).
Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation as defined in Definition C.4, and suppose further that $\sigma(x)=1$ for all $x\in\Sigma^{*}$. Let $p(y\mid x)$ denote the uniform distribution over the solution set $R(x)\coloneqq\{y\mid(x,y)\in R\}$. Then, for CoT, the next-token prediction at each position $i$ corresponds to estimating the relative size of the subproblem
$p(y_{i}\mid y_{<i}=\bar{y}_{<i})\propto\left|R(x^{\prime})\right|,$ (8)
where $x^{\prime}=\psi(x,\bar{y}_{<i}\cdot y_{i})$, i.e., reduces to an instance of the same relation $R$.
We now present the result on the ability of a weak CoT to serve as an FPAUS.
Theorem 5.13 (CoT as Approximate Sampler).
Let $R$ be a self-reducible relation, and let $p(\cdot\mid x)$ denote the uniform distribution over the solution set $R(x)\coloneqq\{y\in\Sigma^{*}\mid(x,y)\in R\}$. There exists a function $\alpha:\mathbb{N}\to\mathbb{R}_{>0}$ such that if a CoT model $\pi$, with vocabulary $\mathcal{V}=\Sigma$, satisfies $(\alpha,\gamma)$-weak token-wise learning with respect to $p$ for some constant $\gamma\in(0,\tfrac{1}{2})$, then there exists a randomized algorithm $\mathcal{A}$ with oracle access to $\pi$ such that, for a given error parameter $\varepsilon>0$, $\mathcal{A}$ outputs a distribution $q^{\pi}_{\varepsilon}(\cdot\mid x)$ satisfying
$\bigl{\|}\,q^{\pi}_{\varepsilon}(\cdot\mid x)\;-\;p(\cdot\mid x)\bigr{\|}_{\mathrm{TV}}\;\;\leq\;\;\varepsilon.$ (9)
Moreover, the total running time of $\mathcal{A}$ is bounded by a polynomial in $(|x|,\log(1/\varepsilon))$.
Remark. The above result applies to any relation for which the CoT satisfies the weak token-wise learning assumption. However, the result becomes computationally meaningful and practically implementable when the relation is self-reducible, as shown in Proposition 5.12.
6 Conclusion
We formally compared CoT and Looped TF, highlighting their respective strengths. Specifically, Looped TF enables efficient parallel computation for deterministic problems, while CoT enables probabilistic inference. Specifically, we introduced a novel perspective grounded in the complexity of approximate computation. These findings offer guidance on whether to CoT or to Loop when scaling inference compute. Future work includes examining the limitations of Looped TF on deterministic tasks, particularly those arising from their fixed space complexity. Another important direction is to compare their FLOPs.
Appendix A Formal Definitions for Section 2
A.1 Transformer Block
We define the computational components of a Transformer using the notation of Merrill & Sabharwal (2025). Let $\mathbb{Q}_{p(n)}$ denote the set of $p(n)$-bit fixed-point numbers with truncated arithmetic.
Definition A.1 (Transformer).
A Transformer consists of the following components:
  1. 1.
    A word embedding function $\mathbf{WE}:{\mathcal{V}}\to\mathbb{Q}_{p}^{m}$, where ${\mathcal{V}}$ denotes the vocabulary set.
  2. 2.
    A time-dependent positional embedding function $\mathbf{PE}:\mathbb{N}\times\mathbb{N}\to\mathbb{Q}_{p}^{m}$.
  3. 3.
    A saturated multi-head self-attention layer $\mathbf{SA}:\mathbb{Q}_{p}^{m\times N}\to\mathbb{Q}_{p}^{m\times N}$ for arbitrary sequence length $N$, parameterized by a matrix $\mathbf{W}:\mathbb{Q}_{p}^{s\times H}\to\mathbb{Q}_{p}^{m}$ and, for each head $h\in[H]$ with head size $s$, matrices ${\mathbf{Q}}^{h},{\mathbf{K}}^{h},{\mathbf{V}}^{h}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{s}$. Given an input $\mathbf{z}_{i}\in\mathbb{Q}_{p}^{m}$ for each position $i\in[n]$, it computes the query $\mathbf{q}_{i,h}={\mathbf{Q}}^{h}\mathbf{z}_{i}$, key $\mathbf{k}_{i,h}={\mathbf{K}}^{h}\mathbf{z}_{i}$, and value $\mathbf{v}_{i,h}={\mathbf{V}}^{h}\mathbf{z}_{i}$, and outputs $\mathbf{W}\cdot\operatorname{concat}(\mathbf{a}_{i,1},\ldots,\mathbf{a}_{i,H}),$ where each attention output $\mathbf{a}_{i,h}$ is defined as:
    $\mathbf{a}_{i,h}=\lim_{\alpha\to\infty}\sum_{j=1}^{c(i)}\frac{\exp(\alpha\,\mathbf{q}_{i,h}^{\top}\mathbf{k}_{j,h})}{Z_{i,h}}\cdot\mathbf{v}_{j,h},\quad Z_{i,h}=\sum_{j=1}^{c(i)}\exp(\alpha\,\mathbf{q}_{i,h}^{\top}\mathbf{k}_{j,h}),$
    with $c(i)=i$ for causal attention and $c(i)=n$ for full attention.
  4. 4.
    A feedforward layer $\mathbf{FF}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{m}$ can take one of the following two forms:
    1. (a)
      The standard layer with parameter matrices $\mathbf{W}_{1}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{w}$ and $\mathbf{W}_{2}:\mathbb{Q}_{p}^{w}\to\mathbb{Q}_{p}^{m}$, where $w$ is the hidden dimension. Given an input $\mathbf{z}_{i}\in\mathbb{Q}_{p}^{m}$, it outputs $\mathbf{W}_{2}\mathrm{ReLU}(\mathbf{W}_{1}\mathbf{z}_{i})$.
    2. (b)
      MoE layer consisting of $N_{E}$ expert $\{\mathbf{W}_{1}^{e},\mathbf{W}_{2}^{e}\}_{e=1}^{N_{E}}$, and a gating matrix $\mathbf{S}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{N_{E}}$. Given an input $\mathbf{z}_{i}\in\mathbb{Q}_{p}^{m}$, it outputs $\mathbf{z}_{i}^{\prime}=\mathbf{W}_{2}^{e^{*}}\,\mathrm{ReLU}(\mathbf{W}_{1}^{e^{*}}\mathbf{z}_{i}),\text{where }e^{*}=\arg\max(\mathbf{S}\mathbf{z}_{i}).$
  5. 5.
    An output function $\bm{\gamma}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{|{\mathcal{V}}|}$, parameterized as a linear transformation.
A.2 Chain-of-Thought
Definition A.2 (Formal: CoT).
Let the Transformer be defined as the composition:
$f\coloneq\bm{\gamma}\circ(\mathsf{id}+\mathbf{FF}_{L})\circ(\mathsf{id}+\mathbf{SA}^{\mathrm{dec}}_{L})\circ\cdots\circ(\mathsf{id}+\mathbf{FF}_{1})\circ(\mathsf{id}+\mathbf{SA}^{\mathrm{dec}}_{1})\circ(\mathbf{WE}+\mathbf{PE}),$
where $\mathbf{SA}^{\mathrm{dec}}_{\ell}$ and $\mathbf{FF}_{\ell}$ denote the causal self-attention and the feedforward layers at depth $\ell\in[L]$, respectively, and $\mathsf{id}$ denotes the identity function. The input tokens are first embedded via the word embedding function $\mathbf{WE}$ and the positional encoding $\mathbf{PE}$, and the final output is produced by a linear projection $\bm{\gamma}$. Given an input sequence $x=(x_{1},\dots,x_{n})\in\mathcal{V}^{n}$, we define the initial sequence as: $f^{0}(x)\coloneq x.$ Then, the CoT model computes recursively as:
$f^{k+1}(x)\coloneq f^{k}(x)\cdot\mathrm{Dec}(f(f^{k}(x))),$
where $\cdot$ denotes concatenation of sequences, and $\mathrm{Dec}(\cdot)$ is a decoding function that maps the output logits to a token in $\mathcal{V}$: in the deterministic case, $\mathrm{Dec}(z)\coloneq\arg\max_{i\in[|\mathcal{V}|]}z_{i}$; in the stochastic case, $\mathrm{Dec}(z)\sim\mathrm{Multinomial}({z_{i}}/{\sum_{j}z_{j}})$, assuming $z_{i}>0$ for all $i$. The final output of the CoT model after $T(n)$ steps is defined as the last $m$ tokens of $f^{T(n)}(x)$, where $m$ denotes the output length.
A.3 Looped Transformer
Definition A.3 (Formal: Looped TF).
Let the Transformer block be defined as the composition:
$f\coloneq(\mathsf{id}+\mathbf{FF}_{L})\circ(\mathsf{id}+\mathbf{SA}_{L})\circ\cdots\circ(\mathsf{id}+\mathbf{FF}_{1})\circ(\mathsf{id}+\mathbf{SA}_{1}),$
where $\mathbf{SA}_{\ell}$ and $\mathbf{FF}_{\ell}$ denote the (non-causal) self-attention and feedforward layers at depth $\ell\in[L]$.
Given an input token sequence $x=(x_{1},\dots,x_{n})\in\mathcal{V}^{n}$, the initial hidden state is: $f^{0}(x)\coloneq\mathbf{WE}(x).$ At each loop iteration $k$, the hidden state is updated by:
$f^{k+1}(x)\coloneq f\left(f^{k}(x)+\mathbf{PE}(\cdot,k)\right),$
where $\mathbf{PE}(i,k)$ denotes a loop-dependent positional encoding at position $i$ and iteration $k$. The final output after $T(n)$ loop iterations is given by $\arg\max_{i\in[|\mathcal{V}|]}\bm{\gamma}\circ f^{T(n)}(x),$ and the model’s prediction is defined as the last $m\leq n$ tokens of this projected sequence, where $m$ denotes the output length.
Appendix B Deferred Proofs for Section 4
B.1 Deferred Definition
This section provides the deferred definitions of the Assumptions 4.24.3, and 4.5.
First, we define the notion of polynomially-efficient approximation, in which a function can be closely approximated by a two-layer ReLU network with polynomial width and logarithmic bit precision, while ensuring consistent separation between distinct outputs. This formulation follows the assumption in Feng et al. (2023), but differs in the input and output domains: whereas their setting considers real-valued functions over continuous domains, we focus on symbolic functions $f_{n}:\Sigma^{\ell(n)}\to\Sigma$ over finite alphabets. Accordingly, we formalize the notion as follows.
Definition B.1 (Polynomially-Efficient Approximation).
Let $f_{n}:\Sigma^{\ell(n)}\to\Sigma$ be a function over input strings of length $\ell(n)$, where $\Sigma=\{s_{1},\dots,s_{|\Sigma|}\}$ is a finite alphabet with fixed enumeration. We say that $f_{n}$ is polynomially-efficiently approximated if there exists a two-layer feedforward network with weight matrices $\mathbf{W}_{1}:\mathbb{Q}_{p(n)}^{\ell(n)\cdot|\Sigma|}\to\mathbb{Q}_{p(n)}^{m},\quad\mathbf{W}_{2}:\mathbb{Q}_{p(n)}^{m}\to\mathbb{Q}_{p(n)}^{|\Sigma|},$ where $m\in\mathrm{poly}(n)$ and $p(n)=\lceil\log_{2}n\rceil$, such that for any input $x\in\Sigma^{\ell(n)}$, the network output
$\hat{y}(x)\coloneq\mathbf{W}_{2}\,\mathrm{ReLU}\big{(}\mathbf{W}_{1}\,\bm{\mathrm{emb}}(x)\big{)}\in\mathbb{Q}_{p(n)}^{|\Sigma|}$ (10)
satisfies the following condition: If $f_{n}(x)=s_{j}$, then there exists a constant $\delta>0$ such that
$\hat{y}_{j}(x)>1-\delta\quad\text{and}\quad\hat{y}_{i}(x)<1-\delta\quad\text{for all }i\neq j.$ (11)
Here, $\bm{\mathrm{emb}}(x)\in\{0,1\}^{\ell(n)\cdot|\Sigma|}$ denotes the concatenation of the one-hot encodings of the input tokens $x=(x_{1},\ldots,x_{\ell(n)})$, where each $x_{i}=s_{k}\in\Sigma$ is represented as $\mathrm{emb}(x_{i})\coloneq\mathbf{e}_{k}\in\{0,1\}^{|\Sigma|}$.
Then, we define a bounded fan-in computation graph as one in which each node receives input from at most a constant number of predecessors.
Definition B.2 (Bounded Fan-In).
The computation graph $G_{n}=(V_{n},E_{n})$ has bounded fan-in: there exists a constant $k\in\mathbb{N}$ such that for all $v\in V_{n}$, the in-degree of $v$ is at most $k$, i.e.,
$\forall v\in V_{n},\quad\deg^{-}(v)\leq k.$ (12)
Lastly, we define the parallel space complexity of a DAG as the maximum number of nodes that must be held in memory simultaneously during layer-wise parallel evaluation, including both the nodes in the current layer and any unresolved dependencies from earlier layers that are required to compute future layers.
Definition B.3 (Parallel Space Complexity).
Let $G=(V,E)$ be a directed acyclic graph. For each node $v\in V$, define its depth by:
$\mathrm{depth}(v)\coloneq\max\left\{\ell\in\mathbb{N}\mid\exists\,u_{0}\to u_{1}\to\dots\to u_{\ell}=v\text{ in }G\right\}.$ (13)
Then the canonical layer decomposition is given by:
$L_{\ell}\coloneq\{v\in V\mid\mathrm{depth}(v)=\ell\}$ (14)
for $\ell=0,1,\dots,D$, where $D\coloneq\max_{v\in V}\mathrm{depth}(v)$ is the depth of $G$.
The parallel space complexity of $G$ is defined as
$S(G)\coloneq\max_{\ell\in\{0,\dots,D\}}\left|L_{\ell}\cup\bigcup_{v\in L_{>\ell}}\left(\mathrm{Pred}(v)\cap\bigcup_{s\leq\ell}L_{s}\right)\right|,$ (15)
where $L_{>\ell}\coloneq\bigcup_{s>\ell}L_{s}$, and $\mathrm{Pred}(v)\coloneq\{u\in V\mid(u\to v)\in E\}$.
B.2 Notation and Orthogonal Vectors
We introduce a method for constructing orthogonal vectors under finite-precision arithmetic and define the associated notation, following the results of Li et al. (2024).
Notation
We follow the notation from Li et al. (2024). For any positive integer $s\in\mathbb{N}^{+}$ and $x\in\{0,1,\dots,2^{s}-1\}$, we denote by $\mathsf{bin}_{s}(x)\in\{0,1\}^{s}$ the standard binary representation of $x$ using $s$ bits, defined such that $x=\sum_{i=1}^{s}2^{i}\cdot(\mathsf{bin}_{s}(x))_{i}.$ We further define the signed binary encoding of $x$, denoted by $\mathsf{sbin}_{s}(x)\in\{-1,1\}^{s}$, as $\mathsf{sbin}_{s}(x)=2\cdot\mathsf{bin}_{s}(x)-(1,\ldots,1).$ Let $x,y\in\mathbb{R}^{s}$ be two vectors of the same length. We define their interleaving, denoted by ${x}^{\frown}{y}\in\mathbb{R}^{2s}$, as follows: $({x}^{\frown}{y})_{2i-1}=x_{i},({x}^{\frown}{y})_{2i}=y_{i}$ for all $i\in[s].$ We denote the rounding operation by $\left[\cdot\right]_{s}\triangleq[\cdot]_{0,s}$. Let $\mathbf{1}_{s}$ denote the all-ones vector of length $s$. We also define rounding operation $\left\langle\cdot,\cdot\right\rangle_{s}$, ×$\times_{s}$, and $\sigma_{s}$ for the inner product, product, and softmax, respectively.
To construct approximately orthogonal vectors under finite-precision inner products, Lemma E.3 of Li et al. (2024) shows that the vectors $q_{i}$ and $k_{j}$ are designed so that the finite-precision inner product $\langle q_{i},k_{j}\rangle_{s}$ vanishes when $i=j$, and accumulates a non-zero bias otherwise. This property plays a crucial role in enabling selective activation in softmax attention.
Lemma B.4 (Li et al., 2024).
For any $s\in\mathbb{N}^{+}$, let $\mathbf{q}_{i}={\mathsf{sbin}_{s}(i)}^{\frown}{1_{s}}$ and $\mathbf{k}_{i}=2^{s+1}\cdot({\mathsf{sbin}_{s}(i)}^{\frown}{(-1_{s})})$ for all $i\in[2^{s}-1]$, it holds that $\left[\exp(\left\langle\mathbf{q}_{i},\mathbf{k}_{j}\right\rangle_{s})\right]_{s}=\mathbf{1}[i=j]$ for all $i,j\in[2^{s}-1]$.
B.3 Proof for Theorem 4.4
Proof.
Our proof follows the general strategy of Theorem 3.3 in Li et al. (2024), which analyzes CoT decoding over Boolean circuits. Here, we extend the argument to general DAG-based computations.
Let $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$ be a directed acyclic computation graph, where each node $v\in V_{n}$ is labeled by a one-hot vector $\bm{\ell}(v)\in\{0,1\}^{|\mathcal{F}_{n}|}$ indicating the function assigned to $v$ from the finite set $\mathcal{F}_{n}$. Let $v_{1},\dots,v_{|V_{n}|}$ denote a fixed topological ordering of $V_{n}$, with inputs appearing first and outputs last. By Assumption 4.3, each node $v_{i}$ has at most $C\in\mathbb{N}$ predecessors. We encode the predecessor indices as a vector $\mathrm{pred}(v_{i})\in[|V_{n}|]^{C}$, with zero-padding where necessary.
Let the vocabulary ${\mathcal{V}}=\Sigma$, and let the finite-precision parameter be $p(n)=\lceil\log_{2}n\rceil$. We assume that the CoT model, implemented as a single-layer decoder-only Transformer, generates the value at each node sequentially following the topological order of the computation graph.
At the decoding step $k$, the model has access to the sequence
$(x_{1},\,x_{2},\,\ldots,\,x_{n},\,y_{1},\,y_{2},\,\ldots,\,y_{k})\in\Sigma^{n+k},$ (16)
where $x=(x_{1},\ldots,x_{n})\in\Sigma^{n}$ denotes the input, and $y_{i}=v_{i}(x)\in\Sigma$ denotes the value computed at node $v_{i}$ for $i\leq k$. We prove by induction that the model can correctly generate $y_{k+1}=v_{k+1}(x)$.
Embedding.
The embedding at position $i$ is a vector in $\mathbb{Q}_{p(n)}^{m}$, where $m\coloneq|{\mathcal{V}}|+|\mathcal{F}_{n}|+(1+C)p(n)+C.$ The embedding is defined as
[…]
where ${\bm{e}}\colon\Sigma\to\mathbb{N}^{|{\mathcal{V}}|}$ is the one-hot encoding, and $\bm{\mathrm{sbinpred}}_{p(n)}(v)\in\mathbb{Q}^{C\cdot p(n)}_{p(n)}$ encodes the binary representations of the predecessor indices:
$\bm{\mathrm{sbinpred}}_{p(n)}(v)=\left(\mathsf{sbin}_{p(n)}(\mathrm{pred}(v)_{0}),\,\ldots,\,\mathsf{sbin}_{p(n)}(\mathrm{pred}(v)_{C-1})\right).$ (18)
This embedding is constructed via the sum of a word embedding $\mathbf{WE}\colon{\mathcal{V}}\to\mathbb{Q}_{p(n)}^{m}$ and a positional embedding $\mathbf{PE}\colon\mathbb{N}\to\mathbb{Q}_{p(n)}^{m}$, given by:
[…]
The final embedding at position $i$ is then $\mathbf{WE}(v_{i}(x))+\mathbf{PE}(i)$.
We then provide the construction of a single-layer decoder-only Transformer.
Attention Layer.
The attention layer consists of $C$ heads. The $h$-th head is configured to attend to the position corresponding to the $h$-th predecessor of node $v_{i+1}$. Specifically, for each position $i\in[n+k]$ and head $h\in[C]$, the attention vectors are defined as:
$\mathbf{q}_{i,h}={\mathsf{sbin}_{p(n)}(\mathrm{pred}(v_{i})_{h})}^{\frown}{1_{p(n)}},\,\mathbf{k}_{i,h}=2^{p(n)+1}\cdot{\mathsf{sbin}_{p(n)}(i)}^{\frown}{(-1_{p})},\,\mathbf{v}_{i,h}=({\bm{e}}(v_{i}(x)),0,\ldots,0).$ (21)
With appropriate output projection, attention retrieves the one-hot representations of each predecessor’s value. The output at position $i$ from the attention layer is:
$\Big{(}{\bm{e}}(v_{i}(x)),\,\bm{\ell}(v_{i+1}),\,\mathsf{sbin}_{p(n)}(i),\,\bm{\mathrm{sbinpred}}_{p(n)}(v_{i+1}),\,\underbrace{{\bm{e}}(v_{\mathrm{pred}(v_{i+1})_{1}}(x)),\,\ldots,\,{\bm{e}}(v_{\mathrm{pred}(v_{i+1})_{C}}(x))}_{\text{updated}}\Big{)}.$ (22)
Feed-forward Layer.
By Assumption 4.2 and Definition B.1, each function $f\in\mathcal{F}_{n}$ can be efficiently approximated by a two-layer neural network. Since $\mathcal{F}_{n}$ is finite, we construct a two-layer mixture-of-experts (MoE) network with one expert per function.
For any node $v$, given the concatenated vector
${\bm{z}}^{(v)}\coloneq\left({\bm{e}}\big{(}v_{\mathrm{pred}(v)_{1}}(x)\big{)},\,\ldots,\,{\bm{e}}\big{(}v_{\mathrm{pred}(v)_{C}}(x)\big{)}\right),$ (23)
the feed-forward output is computed as:
$\widehat{{\bm{e}}(v(x))}=\mathbf{W}_{2}^{(\bm{\ell}(v))}\cdot\mathrm{ReLU}\left(\mathbf{W}_{1}^{(\bm{\ell}(v))}{\bm{z}}^{(v)}\right),$ (24)
where $\mathbf{W}_{1}^{(\bm{\ell}(v))},\mathbf{W}_{2}^{(\bm{\ell}(v))}$ are the parameters of the selected expert. That is, there exists a polynomial-size two-layer feed-forward network which, given input of the form form equation 22, computes an approximation $\widehat{{\bm{e}}(v_{i+1}(x))}\in\mathbb{Q}^{|\Sigma|}_{p(n)}$, where the expert is selected according to $\bm{\ell}(v_{i+1})$.
Decoding.
The next output symbol is then decoded as:
$v_{i+1}(x)=s_{j}\quad\text{where}\quad j=\arg\max_{i\in[|\Sigma|]}\widehat{{\bm{e}}(v_{i+1}(x))}_{i}.$ (25)
This selection can equivalently be expressed as an $\arg\min$ over the negated scores.
Thus, the CoT model correctly computes $y_{k+1}=v_{k+1}(x)$, completing the inductive step. By induction on $k$, the model computes the values at all nodes in topological order, and the proof is complete. ∎
B.4 Proof for Theorem 4.7
Generally, directed acyclic graphs (DAGs) may contain long-range dependencies that span multiple layers, complicating layer-wise simulation by Looped Transformers. To facilitate such simulation, it is beneficial to constrain the dependency structure so that each node depends only on nodes in the immediately preceding layer. To this end, we introduce a transformation called strict layered expansion, which converts any DAG into an equivalent form in which all edges connect adjacent layers by inserting identity-copy nodes that forward intermediate values.
Definition B.5 (Strict Layered Expansion).
Let $G=(V,E,\mathcal{F})$ be a computation graph equipped with a layer decomposition $(L_{0},\dots,L_{D})$, where each node $v\in V$ belongs to some layer $L_{\ell}$ according to its depth $\mathrm{depth}(v)=\ell$. The strict layered expansion of $G$, denoted by $G^{\prime}=(V^{\prime},E^{\prime},\mathcal{F}^{\prime})$, is defined as follows: For each edge $(u\to v)\in E$, where $\mathrm{depth}(u)=d_{u}$ and $\mathrm{depth}(v)=d_{v}$, insert the path
$u^{(d_{u})}\to u^{(d_{u}+1)}\to\cdots\to u^{(d_{v}-1)}\to v^{(d_{v})}$
with each node has identity function, into $E^{\prime}$ by introducing intermediate identity nodes as necessary. Then $G^{\prime}$ is strictly layered: all edges connect adjacent layers, and each dependency is limited to the immediately preceding layer.
Importantly, this transformation preserves both the parallel space complexity (as defined in Definition B.3) and the parallel time complexity (i.e., the depth of the DAG). Then, we show that a Looped Transformer can simulate the strict layered expansion of a computation graph by computing each layer in a single loop iteration.
Proof of Theorem 4.7.
Let $G_{n}^{\prime}=(V_{n}^{\prime},E_{n}^{\prime},\mathcal{F}_{n}^{\prime})$ denote the strict layered expansion of the computation graph $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$. Here, $\mathcal{F}_{n}^{\prime}$ extends $\mathcal{F}_{n}$ by including the identity function, which is assigned to all newly inserted nodes. Define a labeling function $\ell:V_{n}^{\prime}\to\{0,1\}^{|\mathcal{F}_{n}|+1}$ that assigns to each node $v\in V_{n}^{\prime}$ the one-hot encoding of its associated function.
By Assumption 4.5, there exists a constant $W>0$ such that the maximum layer width of $G_{n}^{\prime}$ is at most $Wn$. We fix a consistent ordering within each layer and pad with dummy nodes so that each layer contains exactly $Wn$ nodes. Let $v_{t,i}$ denote the $i$-th node in layer $t$, where $t\in\{0,\dots,D\}$ and $i\in\{1,\dots,Wn\}$, and let $v_{t,i}(x)$ denote the value it computes on input $x$.
By Assumption 4.3, each node $v_{t,i}$ has at most $C\in\mathbb{N}$ predecessors in the previous layer $t-1$. We define $\mathrm{pred}(v_{t,i})\in[Wn]^{C}$ as a vector encoding the indices $j$ such that $v_{t-1,j}\in\mathrm{Pred}(v_{t,i})$, and set unused entries to 00.
We now describe a two-layer Looped Transformer architecture that simulates the layerwise evaluation of the computation graph.
Representation.
Let ${\bm{h}}^{(t,i)}\in\mathbb{Q}_{p(n)}^{mW}$ be the hidden embedding at position $i\in[n]$ after the $t$-th loop iteration, where $p(n)\coloneq\lceil\log_{2}n\rceil$, and $m\coloneq|{\mathcal{V}}|+|\mathcal{F}_{n}|+1+(1+C)p(n)+C.$ Each ${\bm{h}}^{(t,i)}$ is divided into $W$ blocks of size $m$, where the $w$-th block corresponds to node $v_{t+1,\,(i-1)W+w}$, and is defined as:
[…]
with the components defined as:
  • ${\bm{e}}(\cdot)\in\mathbb{N}^{|{\mathcal{V}}|}$: one-hot encoding of the computed value;
  • $\ell(v)\in\mathbb{Q}^{|\mathcal{F}_{n}|+1}_{p(n)}$: one-hot function label;
  • $\mathsf{sbin}_{p(n)}(j)\in\mathbb{Q}^{p(n)}_{p(n)}$: signed binary representation of index $j$;
  • $\bm{\mathrm{sbinpred}}_{p(n)}(v)\in\mathbb{Q}^{C\cdot p(n)}_{p(n)}$: encoding of predecessor indices.
First Layer.
The attention layer consists of $CW$ heads indexed by pairs $(w,c)\in[W]\times[C]$. Each head $h=(w-1)C+c$ retrieves the $c$-th predecessor of node $v_{t+1,(i-1)W+w}$. The query/key/value definitions are:
$\mathbf{q}_{i,h}$ $={\mathsf{sbin}_{p(n)}(\mathrm{pred}(v_{t+1,\,(i-1)W+w})_{c})}^{\frown}{1_{p(n)}},$ (27)
$\mathbf{k}_{i,h}$ $=2^{p(n)+1}\cdot{\mathsf{sbin}_{p(n)}((i-1)W+w)}^{\frown}{-1_{p(n)}},$ (28)
[…]
After aggregation and projection, each block ${\bm{h}}^{(t,i)}_{(w)}$ is updated by writing the retrieved values into the last $C$ slots:
$\begin{split}{\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm}\;\leftarrow\;\\ \Big{(}{\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm-C},\;&\underbrace{e(v_{\mathrm{pred}(v_{t+1,\,(i-1)W+w})_{1}}(x)),\,\ldots,\,e(v_{\mathrm{pred}(v_{t+1,\,(i-1)W+w})_{C}}(x))}_{\text{updated}}\Big{)}.\end{split}$ (30)
Feed-forward Layer.
By Assumption 4.2 and Definition B.1, each $f\in\mathcal{F}_{n}$ is approximable by a two-layer ReLU network. Since $\mathcal{F}_{n}$ is finite, a Mixture-of-Experts (MoE) architecture can be used: for each node $v$, define
${\bm{z}}^{(v)}\coloneq\left({\bm{e}}\big{(}v_{\mathrm{pred}(v)_{1}}(x)\big{)},\,\ldots,\,{\bm{e}}\big{(}v_{\mathrm{pred}(v)_{C}}(x)\big{)}\right),$ (31)
and compute
$\widehat{{\bm{e}}(v(x))}={\bm{W}}_{2}^{(\ell(v))}\cdot\mathrm{ReLU}\left({\bm{W}}_{1}^{(\ell(v))}{\bm{z}}^{(v)}\right).$ (32)
Thus, a polynomial-size two-layer FFN can compute ${\bm{e}}(v_{t+1,\,(i-1)W+w}(x))$ from equation 26. We update the embedding as:
${\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm}\leftarrow\Big{(}\widehat{{\bm{e}}(v_{t+1,\,(i-1)W+w}(x))},\;{\bm{h}}^{(t,i)}_{1+(w-1)m+|{\mathcal{V}}|\,:\,wm}\Big{)}.$ (33)
Second Layer.
The attention layer is disabled (e.g., all attention weights are set to zero). To produce one-hot output, define
$\widehat{{\bm{e}}(f_{n}(x))}_{i}\coloneq\mathrm{ReLU}\left(\hat{y}_{i}(x)-(1-\delta)\right),$ (34)
assuming the correct index $j$ satisfies $\hat{y}_{j}(x)>1-\delta$ and all others are below that threshold. We then update:
${\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm}\leftarrow\Big{(}{\bm{e}}(v_{t+1,\,(i-1)W+w}(x)),\;{\bm{h}}^{(t,i)}_{1+(w-1)m+|{\mathcal{V}}|\,:\,wm}\Big{)}.$ (35)
Positional Embedding.
For the next iteration $t+1$, define the time-dependent positional embedding:
[…]
Thus, the hidden representation after the $(t+1)$-th loop matches the structure in Eq. equation 26. The output function returns the one-hot encoding of the output nodes with negated values, completing the proof. ∎
B.5 Proof for Theorem 4.11
We first present the deferred definition of the class $\mathsf{SIZE}^{\mathcal{L}}$ and state the lemma showing how feedforward layers can simulate gates with unbounded fan-in.
Definition B.6 ($\mathsf{SIZE}^{\mathcal{L}}$ (Li et al., 2024)).
For any decision problem ${\mathcal{L}}$ and $T(n)\subseteq O(\mathsf{poly}(n))$, we define $\mathsf{SIZE}^{\mathcal{L}}(T(n))$ as the set of decision problems ${\mathcal{L}}^{\prime}$ such that there exists $p(n)\in\mathsf{poly}(n)$ and circuits $\{C_{n}\}_{n=1}^{\infty}$ where $C_{n}$ contains at most $O(T(n))$ $\mathsf{AND}$, $\mathsf{OR}$, $\mathsf{NOT}$, and ${\mathcal{L}}_{p(n)}$ gates. For a complexity class ${\mathcal{C}}$, we define $\mathsf{SIZE}^{{\mathcal{C}}}(T(n))\triangleq\cup_{{\mathcal{L}}\in{\mathcal{C}}}\mathsf{SIZE}^{{\mathcal{L}}}(T(n))$.
Lemma B.7 (Li et al., 2024).
Unlimited-fanin $\mathsf{AND},\mathsf{OR}$ (resp. $\mathsf{MAJORITY}):\{0,1\}^{n}\to\{0,1\}$ can be simulated by some 2-layer feedforward ReLU network with constant (resp. $\log{n}$) bits of precision constant hidden dimension and additional $n$ constant inputs of value 1.
Mathematically, let $\mathbf{FF}[s(n)]$ be the set of functions $C:\{0,1\}^{n}\to\{0,1\}$ which can be a two-layer feedforward ReLU network with at most $s(n)$ bits of precision and constant hidden dimension $\mathbf{FF}_{\theta}:\{0,1\}^{2n}\to\{0,1\},\mathbf{FF}_{\theta}(x^{\prime})=\mathbf{W}_{2}\,\mathrm{ReLU}(\mathbf{W}_{1}x^{\prime}+b)$, such that for any $x\in\{0,1\}^{n}$,
$\mathbf{FF}_{\theta}(x_{1},1,x_{2},1,\ldots,x_{n},1)=C(x).$ (37)
We have unlimited-fanin $\mathsf{AND},\mathsf{OR}\in\mathbf{FF}[1]$ and $\mathsf{MAJORITY}\in\mathbf{FF}[\log{n}]$.
We then provide the deferred proof for Looped Transformers.
Proof for Theorem 4.11.
We show that the result for $\mathsf{LOOP}[T(n),\mathrm{poly}(n),\log n]$. The same argument applies to the result for $\mathsf{LOOP}[T(n),\mathrm{poly}(n),1]$. Let $\mathsf{TC}^{[T(n)]}$ denote the class of languages decidable by a nonuniform family of threshold Boolean circuits of polynomial size and depth $T(n)$.
$\mathsf{LOOP}[T(n),\mathsf{poly}(n),\log n]\subseteq\mathsf{TC}^{[T(n)]}$: For each loop iteration, it suffices to show that the update of each bit in each dimension of the embedding at each input position is computable by a $\mathsf{TC}^{0}$ circuit. This follows from the result of Merrill & Sabharwal (2023), which showed that Transformer blocks with logarithmic precision can be simulated by $\mathsf{TC}^{0}$ circuits.
$\mathsf{TC}^{[T(n)]}\subseteq\mathsf{LOOP}[T(n),\mathsf{poly}(n),\log n]$: We show that any threshold circuit of depth $T(n)$ and polynomial size can be simulated by a $\mathsf{LOOP}[T(n),\mathsf{poly}(n),\log n]$ model. We construct a model in which the attention layer, in the first loop iteration, (1) aggregates the input information into a single token position. At each loop iteration, (2) the (looped) feedforward layer excutes the gates corresponding to each layer of the threshold circuit using Lemma B.7.
First, we show how an attention layer combined with a feed-forward layer can be used to copy all input tokens to the final position, i.e.,$\sum_{j=1}^{n}{\bm{e}}_{j}x_{j},$ where ${\bm{e}}_{j}\in\mathbb{N}^{n}$ denotes the one-hot positional vector for index $j$, following the strategy in the proof of Theorem 3.7 in Li et al. (2024).
Each input token at position $i\in[n]$ is embedded as a pair $(x_{i},{\bm{e}}_{i})$. To compute the mapping
$(x_{i},{\bm{e}}_{i})\mapsto(x_{i},x_{i}\cdot{\bm{e}}_{i}),$ (38)
we define a feedforward layer that takes input $(z_{0},z_{1},\dots,z_{n})$ and computes:
$(z_{0},z_{1},\dots,z_{n})\mapsto\left(z_{0},\mathrm{ReLU}(z_{0}+z_{1}-1),\dots,\mathrm{ReLU}(z_{0}+z_{n}-1)\right).$ (39)
To aggregate all positions via uniform attention, we use a single-head attention layer with:
$\mathbf{q}_{i}=\mathbf{k}_{i}=\mathbf{1}_{n},\quad\mathbf{v}_{i}=nx_{i}\cdot{\bm{e}}_{i}\quad\text{for all }i\in[n],$ (40)
where $\mathbf{1}_{n}$ is the all-ones vector in $\mathbb{R}^{n}$, and outputs become $\sum_{i=j}^{n}\frac{1}{n}\cdot\mathbf{v}_{j}=\sum_{i=j}^{n}x_{j}{\bm{e}}_{j}.$
Then, we show that a looped feedforward layer can simulate the circuit. Let the threshold circuit be $G=(V,E)$, where each vertex $v_{i}\in V$ corresponds to a Boolean function $v_{i}(x)$ computed at that node for the input $x$. By Lemma B.7, for each node $v_{i}$, there exists a feedforward layer $\mathbf{FF}_{\theta(i)}$ s.t.
$\mathbf{FF}_{\theta(i)}\left(v_{j_{1}}(x),1,\,v_{j_{2}}(x),1,\,\ldots,\,v_{j_{d_{i}}}(x),1\right)=v_{i}(x),$ (41)
where $j_{1},\dots,j_{d_{i}}$ are the indices of the nodes with edges into $v_{i}$, and $d_{i}=\deg^{-}(v_{i})$.
Then, we define a feedforward layer $\mathbf{FF}_{\theta}$ that simultaneously updates all gate values in parallel. Let $\ell:[|V|]\to V$ be any fixed bijection assigning an index to each gate, and define an embedding vector ${\bm{z}}\in\mathbb{R}^{2|V|}$, where each gate $\ell(i)$ is associated with a 2-dimensional slot $({\bm{z}}_{2i},{\bm{z}}_{2i+1})$. We use ${\bm{z}}_{2i}$ to store the current value $x_{\ell(i)}$, and fix ${\bm{z}}_{2i+1}=1$ as a constant bias input.
The layer $\mathbf{FF}_{\theta}$ internally applies the appropriate local feedforward function $\mathbf{FF}_{\theta(\ell(i))}$ for each gate $\ell(i)$, using the current values of its input gates from the embedding. Formally, for each $i\in[|V|]$, the update is given by
${\bm{z}}_{2i}\leftarrow\mathbf{FF}_{\theta(\ell(i))}\left({\bm{z}}_{2j_{1}},1,{\bm{z}}_{2j_{2}},1,\ldots,{\bm{z}}_{2j_{d_{i}}},1\right),$ (42)
where $\ell(j_{1}),\dots,\ell(j_{d_{i}})$ are the gates with edges into $\ell(i)$.
After $T(n)$ loop iterations—where $T(n)$ is the depth of the circuit DAG—the embedding stores the correct output value $x_{v}$ for every gate $v\in V$.
By appropriately combining the components described above, we construct a $3$-layer Transformer that computes the desired operation. We set the embedding size to be $m=1+n+2|V|$. Given an input sequence $(x_{1},x_{2},\dots,x_{n})\in\{0,1\}^{n}$, each input token $i\in[n]$ is mapped, in the token and positional embedding step, to the initial representation
[…]
where ${\bm{e}}_{i}\in\{0,1\}^{n}$ denotes the one-hot encoding of position $i$, and $\mathbf{0}_{2|V|}\in\mathbb{R}^{2|V|}$ is a zero vector. For each loop iteration $k$, given the input from the previous iteration in the form
$(x_{i},\,{\bm{e}}_{i},\,h_{1}^{(k)},\,h_{2}^{(k)},\,\dots,\,h_{n+2|V|}^{(k)}),$ (44)
the model processes the input and applies computations across layers as follows: In the first two layers, the model extracts and aggregates the input values $x_{1},\dots,x_{n}$. Then, in the third layer, the feedforward layer computes the circuits. We describe the computation for each layer as follows:
  1. 1.
    First layer: The attention layer is disabled (e.g., all attention weights are set to zero), so the input is passed through unchanged. The feed-forward layer computes $x_{i}{\bm{e}}_{i}$ for each position:
    $(x_{i},\,\underbrace{x_{i}{\bm{e}}_{i}}_{\text{updated}},\,\mathbf{0}_{2}n,\,h_{2n+1}^{(k)},\,\dots,\,h_{n+2|V|}^{(k)}),$ (45)
  2. 2.
    Second layer: The attention layer aggregates the $x_{i}{\bm{e}}_{i}$ values into the last token position using uniform attention, as described earlier, producing:
    $(x_{i},\,x_{i}{\bm{e}}_{i},\,\underbrace{x_{1},\,1,\,\dots,\,x_{n},\,1}_{\text{updated}},\,h_{2n+1}^{(k)},\,\dots,\,h_{n+2|V|}^{(k)}).$ (46)
    This ensures that the input vector $x=(x_{1},\dots,x_{n})$ is correctly stored and made available at every subsequent iteration. (The feed-forward layer is disabled.)
  3. 3.
    Third layer: The attention layer is disabled, and the feed-forward layer begins simulating the desired gate computations as defined in Eq. 42, where the input nodes correspond to the first $n$ components of the embedding.
Finally, the output function selects the value of the output gate. This completes the construction. ∎
B.6 Definitions and Proofs for Proposition 4.14
We introduce deferred definitions for each problem and provide proofs showing that each problem lies in nonuniform $\mathsf{NC}^{2}$, thereby proving the statement.
Definition B.8 (Solving Integer Linear Systems).
Given an $n\times n$ integer matrix $A\in\mathbb{Z}^{n\times n}$ and a vector $b\in\mathbb{Z}^{n}$, the task is to determine whether there exists a solution $x\in\mathbb{Q}^{n}$ to the system of linear equations $Ax=b$, and to compute such a solution if it exists.
Proposition B.9.
The problem of solving integer linear systems over $\mathbb{Q}$, where $A\in\mathbb{Z}^{n\times n}$ is non-singular, lies in $\mathsf{NC}^{2}$.
Proof Sketch.
Using Csanky’s algorithm, the inverse of an $n\times n$ matrix over the rationals (or over a field of characteristic zero) can be computed in $\mathsf{NC}^{2}$ (Csanky, 1975). Once $A^{-1}$ is computed, the solution is given by $x=A^{-1}b$, which is a matrix-vector multiplication and hence also lies in $\mathsf{NC}^{1}$. Therefore, the full solution procedure belongs to $\mathsf{NC}^{2}$. ∎
Definition B.10 (Perfect Matching in Bipartite Graphs).
Given a bipartite graph $G=(U\cup V,E)$, where $|U|=|V|=n$, decide whether $G$ contains a perfect matching—that is, a subset of edges $M\subseteq E$ such that every vertex in $U$ and $V$ is incident to exactly one edge in $M$.
Proposition B.11.
The bipartite perfect matching problem is in nonuniform $\mathsf{NC}^{2}$.
Proof Sketch.
The problem can be reduced to checking whether a certain symbolic matrix has a non-zero determinant, and the determinant of an $n\times n$ matrix can be computed in randomized uniform $\mathsf{NC}^{2}$ (Karp et al., 1985). By Adleman’s theorem (Adleman, 1978), any $\mathsf{BPP}$ algorithm can be simulated in $\mathsf{P}/\mathsf{poly}$, meaning that for each input length $n$, the randomness used by the algorithm can be fixed in advance and hard-wired as advice. This yields a nonuniform circuit family of depth $O(\log^{2}n)$ and polynomial size. Thus, the problem lies in nonuniform $\mathsf{NC}^{2}$. ∎
Definition B.12 (Fixed Context-Free Language Recognition).
Let $G$ be a fixed context-free grammar. Given an input string $w\in\Sigma^{*}$, decide whether $w\in L(G)$, i.e., whether $w$ is generated by $G$.
Note that this is the nonuniform word problem, and is distinct from the uniform CFG membership problem, where the grammar is part of the input and the problem is known to be $\mathsf{P}$-complete.
Proposition B.13 (Ruzzo, 1979).
The word problem for a fixed context-free language is in $\mathsf{NC}^{2}$.
Appendix C Background on Approximate Counting and Sampling
This section provides an overview of key definitions, notation, and foundational results concerning approximate counting and almost uniform generation, with a particular focus on self-reducible relations.
C.1 Definitions of Approximate Counting
Many computational problems can be formalized as counting the number of solutions satisfying a certain predicate. Formally, a counting problem is defined for relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$, where $\Sigma$ is a finite alphabet. For each input $x\in\Sigma^{*}$, the goal is to compute the number of feasible solutions $y\in\Sigma^{*}$ such that $(x,y)\in R$; that is, counting function #$\#R:\Sigma^{*}\to\mathbb{N}$ is defined by #$\#R(x)\coloneq\left|\left\{y\in\Sigma^{*}\mid(x,y)\in R\right\}\right|.$
The complexity class #$\#\mathsf{P}$, introduced by Valiant (1979), captures the class of functions that count the number of accepting paths of a nondeterministic polynomial-time Turing machine.
Definition C.1 (#$\#\mathsf{P}$).
A function $f:\Sigma^{*}\to\mathbb{N}$ belongs to the class #$\#\mathsf{P}$ if there exists a polynomial-time Turing machine $M$ and a polynomial $p$ such that, for every input $x\in\Sigma^{*}$,
$f(x)=\left|\left\{y\in\{0,1\}^{p(|x|)}\mid M(x,y)=1\right\}\right|.$ (47)
Intuitively, #$\#\mathsf{P}$ encompasses problems such as counting the number of satisfying assignments of a Boolean formula (#$\#\mathrm{SAT}$), computing the permanent of a 0,1 matrix, or counting the number of perfect matchings in a bipartite graph (Valiant, 1979). All of these are known to be #$\#\mathsf{P}$-complete.
Approximate Counting
Since exact computation of #$\#\mathsf{P}$ functions is believed to be intractable, much attention has been devoted to the design of efficient approximation algorithms. This motivates the study of fully polynomial randomized approximation schemes (FPRAS) and their connections to approximate uniform generation, which we describe in the subsequent sections.
Definition C.2 (FPTAS).
Let $R$ be a relation, and define $f(x)\coloneq\lvert R(x)\rvert$. A fully polynomial-time approximation scheme (FPTAS) is a deterministic algorithm $A(x,\varepsilon)$ that, for any input $x\in\{0,1\}^{*}$ and any error tolerance $\varepsilon>0$, produces a value $\hat{f}(x)\in\mathbb{Q}$ such that
$(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$ (48)
and the running time of $A$ is bounded by a polynomial in both the input size $|x|$ and $1/\varepsilon$; i.e., $\mathrm{poly}(|x|,1/\varepsilon)$. Let the class $\mathsf{FPTAS}$ denote the set of all functions that admit such a scheme.
Randomized Schemes for Approximation
While an FPTAS provides a deterministic polynomial-time algorithm for approximating counting functions within any desired relative error, such schemes are known to exist only for a very restricted class of problems. The concept of randomized approximation schemes was introduced by Karp & Luby (1983), who proposed using Monte Carlo algorithms to approximate the value of hard counting functions. Their work laid the foundation for what is now known as the fully polynomial randomized approximation scheme (FPRAS), which allows one to efficiently obtain high-probability multiplicative approximations for a broad class of #$\#\mathsf{P}$ functions.
Definition C.3 (FPRAS).
Let $R$ be a relation, and define $f(x)\coloneq\lvert R(x)\rvert$. A randomized approximation scheme is an algorithm that, given any error tolerance $\varepsilon>0$ and failure probability $\delta>0$, outputs an estimate $\hat{f}(x)$ such that
$\Pr\left[(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x)\right]\geq 1-\delta.$ (49)
If the algorithm runs in time polynomial in $|x|$, $1/\varepsilon$, and $\log(1/\delta)$, it is called a Fully Polynomial-Time Randomized Approximation Scheme (FPRAS) for $f$. We denote by $\mathsf{FPRAS}$ the class of functions that admit such an algorithm.
C.2 Complexity of Self-Reducible Problems
The concept of self-reducibility, introduced by Schnorr (1976), plays a central role in bridging approximate counting and approximate sampling. Informally, a relation $R(x)$ is self-reducible if any solution $y\in R(x)$ can be obtained by recursively solving a polynomial number of smaller subproblems derived from $x$.
Definition C.4 (Self-Reducibility (Schnorr, 1976)).
A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is self-reducible if:
  1. 1.
    There exists a polynomial-time computable function $g\in\Sigma^{*}\to\mathbb{N}$ such that $xRy\Rightarrow|y|=g(x);$
  2. 2.
    There exist polynomial-time computable functions $\psi\in\Sigma^{*}\times\Sigma^{*}\to\Sigma^{*}$ and $\sigma\in\Sigma^{*}\to\mathbb{N}$ s.t.
    $\sigma(x)$ $=O(\log|x|),$ (50)
    $g(x)>0$ $\Rightarrow\sigma(x)>0\quad\forall x\in\Sigma^{*},$ (51)
    $|\psi(x,w)|$ $\leq|x|\quad\forall x,w\in\Sigma^{*},$ (52)
    and such that, for all $x\in\Sigma^{*}$, $y=y_{1}\dots y_{n}\in\Sigma^{*}$,
    $\langle x,y_{1},\dots,y_{n}\rangle\in R\iff\langle\psi(x,y_{1}\dots y_{\sigma(x)}),y_{\sigma(x)+1},\dots,y_{n}\rangle\in R.$ (53)
Self-reducibility relation bridges the gap between approximate counting and almost uniform generation. A fundamental result by Jerrum et al. (1986) shows that, under the self-reducibility condition, these two problems have the same time complexity: the ability to approximately count solutions within a polynomial factor implies the ability to sample from an almost uniform distribution over the solution set, and vice versa.
Theorem C.5 (Jerrum et al., 1986).
Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation. There exists an FPRAS for approximating $|R(x)|$ if and only if there exists an FPAUS for sampling uniformly from $R(x)$.
Appendix D Deferred Proofs for Section 5
D.1 Proof for Theorem 5.6
Proof.
By Lemma 5.5, a probabilistic CoT model can simulate a polynomial-time probabilistic Turing machine, and hence can implement any FPRAS algorithm within polynomial CoT steps. In contrast, deterministic Looped Transformers with a polynomial number of loop iterations correspond to deterministic polynomial-time Turing machines, and hence can only compute functions in $\mathsf{FPTAS}$, assuming that each loop iteration performs a fixed computation deterministically.
Given the assumption that $\mathsf{FPTAS}\neq\mathsf{FPRAS}$, there exists a relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ such that approximating $|R(x)|$ to within a relative error $\varepsilon>0$ lies in $\mathsf{FPRAS}\setminus\mathsf{FPTAS}$. Therefore, a probabilistic CoT model can approximate $|R(x)|$ within arbitrary relative error with high probability in polynomial time, whereas no deterministic Looped Transformer with polynomial loop iterations can achieve the same approximation. ∎
D.2 Proof for Lemma 5.11
Proof.
Fix any token position $i\in[m]$. By the definition of $(\alpha,\gamma)$-weak token-wise learning, we have
$\Pr_{r_{i}\sim\pi(r_{i}\mid x,y_{<i})}\left[\mathcal{L}_{\mathrm{CE}}(x,y_{<i},r_{i})\leq\alpha(|x|)\right]\geq\tfrac{1}{2}+\gamma.$ (54)
Since $\mathrm{KL}(p\,\|\,\pi)\leq\mathcal{L}_{\mathrm{CE}},$ it follows that
$\Pr_{r_{i}}\left[\mathrm{KL}\left(p(\cdot\mid x,y_{<i})\,\|\,\pi(\cdot\mid x,y_{<i},r_{i})\right)\leq\alpha(|x|)\right]\geq\tfrac{1}{2}+\gamma.$ (55)
By Pinsker’s inequality, this implies
$\Pr_{r_{i}}\left[\|\pi^{(t)}-p\|_{\mathrm{TV}}\leq\sqrt{\tfrac{1}{2}\alpha(|x|)}\right]\geq\tfrac{1}{2}+\gamma.$ (56)
From the definition of total variation distance,
$\|\pi^{(t)}-p\|_{\mathrm{TV}}=\max_{A\subseteq\mathcal{V}}\left|\pi^{(t)}(A)-p(A)\right|\geq\left|\pi^{(t)}(y)-p(y)\right|\quad\text{for all }y\in\mathcal{V}.$
Hence,
$\Pr_{r_{i}}\left[\left|\pi^{(t)}(y)-p(y)\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\right]\geq\tfrac{1}{2}+\gamma.$ (57)
Let $Z_{t}=\mathbf{1}\left[\left|\pi^{(t)}(y)-p(y)\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\right]$, and note that $\mathbb{E}[Z_{t}]\geq\tfrac{1}{2}+\gamma$. Let $S=\sum_{t=1}^{T}Z_{t}$ denote the number of good samples. Then the expectation satisfies $\mathbb{E}[S]=\sum_{t=1}^{T}\mathbb{E}[Z_{t}]\geq T\left(\tfrac{1}{2}+\gamma\right).$
By a standard Chernoff bound:
$\Pr\left[S<(1-\varepsilon)\mathbb{E}[S]\right]\leq\exp\left(-\frac{\varepsilon^{2}\mathbb{E}[S]}{2}\right).$ (58)
To bound the probability that fewer than half the samples are good, we set
$(1-\varepsilon)\mathbb{E}[S]=\tfrac{T}{2}\quad\Rightarrow\quad\varepsilon=1-\frac{T/2}{\mathbb{E}[S]}\geq\frac{2\gamma}{1+2\gamma}\geq\gamma,$ (59)
where the last inequality holds for $\gamma<1/2$.
Substituting into the Chernoff bound gives
$\Pr\left[\sum_{t=1}^{T}Z_{t}<\tfrac{T}{2}\right]\leq\exp\left(-\frac{\gamma^{2}\cdot\mathbb{E}[S]}{2}\right)\leq\exp(-2T\gamma^{2}),$ (60)
where the final inequality uses $\mathbb{E}[S]\geq T(\tfrac{1}{2}+\gamma)$.
Therefore, if $T\geq\frac{2}{\gamma^{2}}\log\left(\frac{|\mathcal{V}|}{\delta}\right)$, then with probability at least $1-\delta/|\mathcal{V}|$, more than half of the values $\pi^{(t)}(y)$ lie within $O\left(\sqrt{\alpha(|x|)}\right)$ of $p(y)$. Thus, the median $\pi^{\star}(y)$ also lies within this range.
Applying a union bound over all $y\in\mathcal{V}$, we conclude that with probability at least $1-\delta$,
$\left|\pi^{\star}(y)-p(y)\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\quad\text{for all }y\in\mathcal{V}.$ (61)
D.3 Proof for Theorem 5.13
Proof.
The proof follows the approach of rejection sampling based on approximate next-token predictions, inspired by Jerrum et al. (1986).
Let $R(x)\coloneqq\{y\in\Sigma^{*}\mid(x,y)\in R\}$ denote the solution set for input $x\in\Sigma^{*}$, and let $p(\cdot\mid x)$ be the uniform distribution over $R(x)$. By the self-reducibility of $R$, there exists a polynomially bounded function $m:\mathbb{N}\to\mathbb{N}$ such that all $y\in R(x)$ satisfy $|y|\leq m(|x|)$.
Assume the CoT model $\pi$ satisfies $(\alpha,\gamma)$-weak token-wise learning with respect to $p$, for some constant $\gamma\in(0,\tfrac{1}{2})$ and function $\alpha:\mathbb{N}\to\mathbb{R}_{>0}$. Then for each token position $i\in[m(|x|)]$, with probability at least $1/2+\gamma$, the prediction $\pi(y_{i}\mid x,y_{<i})$ approximates the true conditional probability $p(y_{i}\mid x,y_{<i})$ within total variation distance $\leq 1/\alpha(|x|)$. By Pinsker’s inequality, this implies the pointwise deviation:
$\left|\pi(y_{i}\mid x,y_{<i})-p(y_{i}\mid x,y_{<i})\right|\leq\sqrt{\frac{1}{2\alpha(|x|)}}.$ (62)
Since $p(y_{i}\mid x,y_{<i})\geq 1/|R(x)|$, we have the relative error bound:
$\left|\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}-1\right|\leq\sqrt{\frac{1}{2\alpha(|x|)}}\cdot|R(x)|.$ (63)
Choose $\alpha(|x|)\geq\frac{1}{4\cdot m(|x|)^{2}\cdot|R(x)|^{2}}$, so that:
$\left|\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}-1\right|\leq\frac{1}{2m(|x|)}.$ (64)
Then for all $y\in R(x)$, we have:
$\frac{\pi(y\mid x)}{p(y\mid x)}=\prod_{i=1}^{m(|x|)}\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}\in\left[\left(1-\frac{1}{2m(|x|)}\right)^{m(|x|)},\left(1+\frac{1}{2m(|x|)}\right)^{m(|x|)}\right].$ (65)
Using the inequality $\left(1+\tfrac{1}{2m}\right)^{m}\leq e^{1/2}$, we obtain:
$\frac{1}{\beta}\leq\frac{\pi(y\mid x)}{p(y\mid x)}\leq\beta,\quad\text{where }\beta=e^{1/2}.$ (66)
Now define the rejection sampling algorithm:
  • Sample $y\sim\pi(\cdot\mid x)$;
  • If $y\notin R(x)$, reject and repeat;
  • If $y\in R(x)$, accept with probability $\frac{1}{\beta}\cdot\frac{p(y\mid x)}{\pi(y\mid x)}\leq 1$;
This produces exact samples from $p(\cdot\mid x)$. For any $y\in R(x)$, the probability of acceptance is:
$\pi(y\mid x)\cdot\left(\frac{1}{\beta}\cdot\frac{p(y\mid x)}{\pi(y\mid x)}\right)=\frac{1}{\beta|R(x)|},$ (67)
which is independent of $y$. Hence, the output is uniform over $R(x)$.
Error Analysis.
There are two sources of error:
  1. 1.
    The condition in Eq. 62 may fail to hold at some positions $i\in[m(|x|)]$.
  2. 2.
    The rejection sampling procedure may fail to produce any accepted sample within a bounded number of trials.
To address (1), we use the Lemma 5.11 and draw $T=O\left(\frac{1}{\gamma^{2}}\log\left(\frac{|\mathcal{V}|\cdot m(|x|)}{\varepsilon/2}\right)\right)$ independent reasoning paths to estimate each conditional probability. By a union bound over all positions and tokens, the probability that any of them exceeds the deviation bound is at most $\varepsilon/2$.
To address (2), note that the acceptance probability is at least $1/\beta^{2}$. Hence, by running $N=O\left(\beta^{2}\log\left(\frac{2}{\varepsilon}\right)\right)$ independent trials, the probability that all of them are rejected is at most
$(1-1/\beta^{2})^{N}\leq e^{-N/\beta^{2}}\leq\varepsilon/2.$ (68)
Combining both sources of error, with probability at least $1-\varepsilon$, the algorithm outputs a sample from a distribution $q^{\pi}_{\varepsilon}(\cdot\mid x)$ satisfying
$\bigl{\|}q^{\pi}_{\varepsilon}(\cdot\mid x)-p(\cdot\mid x)\bigr{\|}_{\mathrm{TV}}\leq\varepsilon.$ (69)
Running Time.
Each token prediction and rejection trial takes polynomial time in $|x|$, and the number of trials is bounded by $O(\log(1/\varepsilon))$ with high probability. Therefore, the total running time is polynomial in $(|x|,\log(1/\varepsilon))$. ∎

### v2

@@
 These separations suggest the tasks for which depth-driven recursion is more suitable, thereby offering practical cues for choosing between reasoning paradigms.
+This paper is outdated.
+For the latest version, please visit arXiv:2509.25239.
 1 Introduction
@@
 While these results suggest a superiority of looped models over CoT, several important questions remain unanswered:
-  • Does separation exist beyond the logarithmic regime? → Affirmative; see Section 4.
-  • Are Looped Transformers universally more expressive than CoT? → Negative; see Section 5.
+  • Does separation exist beyond the logarithmic regime? → Affirmative; see Section˜4.
+  • Are Looped Transformers universally more expressive than CoT? → Negative; see Section˜5.
 In this paper, we aim to clarify the distinct characteristics of Looped TF and CoT by analyzing two problem settings, thereby addressing these questions.
@@
 For the feedforward computation, we consider either a standard fully connected layer or a mixture-of-experts (MoE) layer (Shazeer et al., 2017; Csordás et al., 2024).
-Formal definitions of the Transformer block are deferred to Appendix A.
+Formal definitions of the Transformer block are deferred to LABEL:app:tf.
 2.1 Chain-of-Thought
@@
 2.3 Notation
-For any $n\in\mathbb{N}^{+}$, we write $[n]\coloneq\{1,2,\ldots,n\}$.
+For any $n\in\mathbb{N}^{+}$, we write $n]\coloneq\{1,2,\ldots,n\}$.
 For functions $f,g:\mathbb{N}\to\mathbb{N}$, we write $f(n)\in O(g(n))$ iff there exist constants $c>0$ and $n_{0}\in\mathbb{N}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$.
@@
 3 General Problem Setting
-This section begins by reviewing the problem settings considered in prior work, as summarized in Table 1, and identifying the remaining open questions.
+This section begins by reviewing the problem settings considered in prior work, as summarized in Table˜1, and identifying the remaining open questions.
 We then formalize the two problem settings addressed in this study, specifying their underlying assumptions and objectives.
@@
 3.3 Our Problem Settings
-We consider two fundamental classes of computational problems: the deterministic problem and the probabilistic problem, elaborated in Section 4 and Section 5, respectively.
+We consider two fundamental classes of computational problems: the deterministic problem and the probabilistic problem, elaborated in Section˜4 and Section˜5, respectively.
 This section provides high-level definitions that specify the input–output structure in each setting; precise formulations are deferred to the corresponding sections.
@@
 This section analyzes the fundamental computational differences between CoT and Looped TF in the deterministic setting.
-We formalize deterministic computations using computation graphs, as introduced in Section 4.1.
-In Section 4.2, we show how CoT and Looped TF differ in processing the same computation graphs, highlighting their respective strengths.
-Finally, Section 4.3 presents a separation in their computational complexity in the polylogarithmic regime.
+We formalize deterministic computations using computation graphs, as introduced in Section˜4.1.
+In Section˜4.2, we show how CoT and Looped TF differ in processing the same computation graphs, highlighting their respective strengths.
+Finally, Section˜4.3 presents a separation in their computational complexity in the polylogarithmic regime.
 4.1 Problem Setting
@@
 A computation graph is a directed acyclic graph $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$ that defines a function $G_{n}:\Sigma^{n}\to\Sigma^{m}$, where $\mathcal{F}_{n}=\{f_{n}:\Sigma^{\ell_{n}}\to\Sigma\mid\ell_{n}\in\mathbb{N}\}$ is a finite set of local functions associated with each node.
-The graph $G_{n}$ consists of (i) input nodes with in-degree 00, (ii) function nodes labeled by $f\in\mathcal{F}_{n}$ whose in-degree matches the arity of the corresponding function, and (iii) $m$ output nodes $O\subseteq V_{n}$ with out-degree 00.
+The graph $G_{n}$ consists of (i) input nodes with in-degree 0, (ii) function nodes labeled by $f\in\mathcal{F}_{n}$ whose in-degree matches the arity of the corresponding function, and (iii) $m$ output nodes $O\subseteq V_{n}$ with out-degree 0.
 The overall function is obtained by evaluating the graph in topological order.
@@
 To characterize the expressiveness of CoT and Looped TF, we analyze how each simulates computation graphs.
-We begin by formalizing the assumptions under which computations can be implemented by Transformer-based models (see Section B.1 for details).
+We begin by formalizing the assumptions under which computations can be implemented by Transformer-based models (see LABEL:app:def for details).
 Assumption 4.2 (Polynomially-Efficient Approximation (cf. Feng et al., 2023)).
@@
 Remark.
-Proofs are given in Appendix B, with illustrations in Fig.
+Proofs are given in LABEL:app:det, with illustrations in Fig.
 1.
-Informally, Theorem 4.4 shows that CoT simulates computations by decoding nodes in topological order, whereas Theorem 4.7 shows that Looped TF simulates computations in a layer-wise and parallel fashion.
+Informally, Theorem˜4.4 shows that CoT simulates computations by decoding nodes in topological order, whereas Theorem˜4.7 shows that Looped TF simulates computations in a layer-wise and parallel fashion.
 This contrast reflects complementary strengths: CoT employs intermediate steps as scratchpad memory, supporting general computation without structural constraints.
@@
 Let $\mathsf{CoT}[T(n),d(n),p(n)]$ (resp.
-$\mathsf{LOOP}[T(n),d(n),p(n)]$) denote the set of languages :${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a CoT (resp.
+$\mathsf{LOOP}[T(n),d(n),p(n)]$) denote the set of languages ${\mathcal{L}}:\{0,1\}^{*}\to\{0,1\}$ for which there exists a CoT (resp.
 Looped TF) $M_{n}$ for each input size $n$, with embedding size $O(d(n))$ and $O(p(n))$ bits of precision, such that for all $x\in\{0,1\}^{n}$, the final output token, after $O(T(n))$ decode steps (resp.
@@
 Prior work has shown that CoT is equivalent to Boolean circuits composed of $\mathsf{TC}^{0}$ gates, with circuit “size” proportional to the number of CoT steps.
-This class $\mathsf{SIZE}^{\mathsf{TC}^{0}}$ is defined in Definition B.6.
+This class $\mathsf{SIZE}^{\mathsf{TC}^{0}}$ is defined in LABEL:def:tcsize.
 Theorem 4.10 (Li et al., 2024).
@@
 Instance Examples
-There are several concrete problems in (nonuniform) $\mathsf{NC}^{2}$: solving linear equations, testing for perfect matching in bipartite graphs, and membership testing for fixed context-free grammars, denoted by LIN, MATCH, and FCFL, respectively (see Section B.6 for details).
+There are several concrete problems in (nonuniform) $\mathsf{NC}^{2}$: solving linear equations, testing for perfect matching in bipartite graphs, and membership testing for fixed context-free grammars, denoted by LIN, MATCH, and FCFL, respectively (see LABEL:app:nc2 for details).
 Then we have:
@@
 In this section, we investigate the expressiveness of CoT with stochastic decoding for probabilistic tasks.
-Section 5.1 introduces the task formulation and defines the model.
-Section 5.2 analyzes the expressive power of CoT in an idealized setting and establishes a formal separation from Looped TF.
-Finally, Section 5.3 shows that even weak CoT, when combined with self-consistency, exhibits sufficient expressive power.
+Section˜5.1 introduces the task formulation and defines the model.
+Section˜5.2 analyzes the expressive power of CoT in an idealized setting and establishes a formal separation from Looped TF.
+Finally, Section˜5.3 shows that even weak CoT, when combined with self-consistency, exhibits sufficient expressive power.
 5.1 Problem Setting
@@
 A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is self-reducible if any solution $y=y_{1}\dots y_{n}$ to input $x$ can be constructed incrementally, such that a prefix $y_{1:\sigma(x)}$ determines a subproblem $\psi(x,y_{1:\sigma(x)})$ whose solution completes $y$.
-See Definition C.4 for a formal definition.
+See LABEL:def:self-red for a formal definition.
 For example, consider the Boolean satisfiability problem (SAT).
@@
 A Fully Polynomial Randomized Approximation Scheme (FPRAS) is a randomized version that additionally takes $\delta>0$, and satisfies the same guarantee with probability at least $1-\delta$.
-See Appendix C for details.
+See LABEL:app:approx for details.
 5.2 CoT is More Expressive than Looped TF in Approximate Inference
@@
 5.3 Expressiveness of Weak CoT
-While Lemma 5.5 analyzes an idealized setting where a CoT, viewed as a PTM, is assumed to receive explicit error parameters, such assumptions diverge from the behavior of practical LLMs.
+While Lemma˜5.5 analyzes an idealized setting where a CoT, viewed as a PTM, is assumed to receive explicit error parameters, such assumptions diverge from the behavior of practical LLMs.
 In practice, models are trained within bounded error, but do not offer explicit control over the accuracy of their outputs.
@@
   • Can a weak CoT be transformed into an approximate sampler for any desired accuracy?
-These questions are addressed in the subsequent sections, with formal proofs deferred to Appendix D and key ideas shown in Fig. 2.
+These questions are addressed in the subsequent sections, with formal proofs deferred to LABEL:app:prob and key ideas shown in Fig. 2.
 Figure 2: Illustration of next-token prediction in CoT as approximate relative counting over a self-reducible relation.
@@
 Proposition 5.12 (Reducing Next-Token Prediction to Relative Counting via Self-Reducibility).
-Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation as defined in Definition C.4, and suppose further that $\sigma(x)=1$ for all $x\in\Sigma^{*}$.
+Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation as defined in LABEL:def:self-red, and suppose further that $\sigma(x)=1$ for all $x\in\Sigma^{*}$.
 Let $p(y\mid x)$ denote the uniform distribution over the solution set $R(x)\coloneqq\{y\mid(x,y)\in R\}$.
@@
 The above result applies to any relation for which the CoT satisfies the weak token-wise learning assumption.
-However, the result becomes computationally meaningful and practically implementable when the relation is self-reducible, as shown in Proposition 5.12.
+However, the result becomes computationally meaningful and practically implementable when the relation is self-reducible, as shown in Proposition˜5.12.
 6 Conclusion
@@
 Another important direction is to compare their FLOPs.
-Appendix A Formal Definitions for Section 2
-A.1 Transformer Block
-We define the computational components of a Transformer using the notation of Merrill & Sabharwal (2025).
-Let $\mathbb{Q}_{p(n)}$ denote the set of $p(n)$-bit fixed-point numbers with truncated arithmetic.
-Definition A.1 (Transformer).
-A Transformer consists of the following components:
-  1. 1.
-    A word embedding function $\mathbf{WE}:{\mathcal{V}}\to\mathbb{Q}_{p}^{m}$, where ${\mathcal{V}}$ denotes the vocabulary set.
-  2. 2.
-    A time-dependent positional embedding function $\mathbf{PE}:\mathbb{N}\times\mathbb{N}\to\mathbb{Q}_{p}^{m}$.
-  3. 3.
-    A saturated multi-head self-attention layer $\mathbf{SA}:\mathbb{Q}_{p}^{m\times N}\to\mathbb{Q}_{p}^{m\times N}$ for arbitrary sequence length $N$, parameterized by a matrix $\mathbf{W}:\mathbb{Q}_{p}^{s\times H}\to\mathbb{Q}_{p}^{m}$ and, for each head $h\in[H]$ with head size $s$, matrices ${\mathbf{Q}}^{h},{\mathbf{K}}^{h},{\mathbf{V}}^{h}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{s}$.
-Given an input $\mathbf{z}_{i}\in\mathbb{Q}_{p}^{m}$ for each position $i\in[n]$, it computes the query $\mathbf{q}_{i,h}={\mathbf{Q}}^{h}\mathbf{z}_{i}$, key $\mathbf{k}_{i,h}={\mathbf{K}}^{h}\mathbf{z}_{i}$, and value $\mathbf{v}_{i,h}={\mathbf{V}}^{h}\mathbf{z}_{i}$, and outputs $\mathbf{W}\cdot\operatorname{concat}(\mathbf{a}_{i,1},\ldots,\mathbf{a}_{i,H}),$ where each attention output $\mathbf{a}_{i,h}$ is defined as:
-    $\mathbf{a}_{i,h}=\lim_{\alpha\to\infty}\sum_{j=1}^{c(i)}\frac{\exp(\alpha\,\mathbf{q}_{i,h}^{\top}\mathbf{k}_{j,h})}{Z_{i,h}}\cdot\mathbf{v}_{j,h},\quad Z_{i,h}=\sum_{j=1}^{c(i)}\exp(\alpha\,\mathbf{q}_{i,h}^{\top}\mathbf{k}_{j,h}),$
-    with $c(i)=i$ for causal attention and $c(i)=n$ for full attention.
-  4. 4.
-    A feedforward layer $\mathbf{FF}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{m}$ can take one of the following two forms:
-    1. (a)
-      The standard layer with parameter matrices $\mathbf{W}_{1}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{w}$ and $\mathbf{W}_{2}:\mathbb{Q}_{p}^{w}\to\mathbb{Q}_{p}^{m}$, where $w$ is the hidden dimension.
-Given an input $\mathbf{z}_{i}\in\mathbb{Q}_{p}^{m}$, it outputs $\mathbf{W}_{2}\mathrm{ReLU}(\mathbf{W}_{1}\mathbf{z}_{i})$.
-    2. (b)
-      MoE layer consisting of $N_{E}$ expert $\{\mathbf{W}_{1}^{e},\mathbf{W}_{2}^{e}\}_{e=1}^{N_{E}}$, and a gating matrix $\mathbf{S}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{N_{E}}$.
-Given an input $\mathbf{z}_{i}\in\mathbb{Q}_{p}^{m}$, it outputs $\mathbf{z}_{i}^{\prime}=\mathbf{W}_{2}^{e^{*}}\,\mathrm{ReLU}(\mathbf{W}_{1}^{e^{*}}\mathbf{z}_{i}),\text{where }e^{*}=\arg\max(\mathbf{S}\mathbf{z}_{i}).$
-  5. 5.
-    An output function $\bm{\gamma}:\mathbb{Q}_{p}^{m}\to\mathbb{Q}_{p}^{|{\mathcal{V}}|}$, parameterized as a linear transformation.
-A.2 Chain-of-Thought
-Definition A.2 (Formal: CoT).
-Let the Transformer be defined as the composition:
-$f\coloneq\bm{\gamma}\circ(\mathsf{id}+\mathbf{FF}_{L})\circ(\mathsf{id}+\mathbf{SA}^{\mathrm{dec}}_{L})\circ\cdots\circ(\mathsf{id}+\mathbf{FF}_{1})\circ(\mathsf{id}+\mathbf{SA}^{\mathrm{dec}}_{1})\circ(\mathbf{WE}+\mathbf{PE}),$
-where $\mathbf{SA}^{\mathrm{dec}}_{\ell}$ and $\mathbf{FF}_{\ell}$ denote the causal self-attention and the feedforward layers at depth $\ell\in[L]$, respectively, and $\mathsf{id}$ denotes the identity function.
-The input tokens are first embedded via the word embedding function $\mathbf{WE}$ and the positional encoding $\mathbf{PE}$, and the final output is produced by a linear projection $\bm{\gamma}$.
-Given an input sequence $x=(x_{1},\dots,x_{n})\in\mathcal{V}^{n}$, we define the initial sequence as: $f^{0}(x)\coloneq x.$ Then, the CoT model computes recursively as:
-$f^{k+1}(x)\coloneq f^{k}(x)\cdot\mathrm{Dec}(f(f^{k}(x))),$
-where $\cdot$ denotes concatenation of sequences, and $\mathrm{Dec}(\cdot)$ is a decoding function that maps the output logits to a token in $\mathcal{V}$: in the deterministic case, $\mathrm{Dec}(z)\coloneq\arg\max_{i\in[|\mathcal{V}|]}z_{i}$; in the stochastic case, $\mathrm{Dec}(z)\sim\mathrm{Multinomial}({z_{i}}/{\sum_{j}z_{j}})$, assuming $z_{i}>0$ for all $i$.
-The final output of the CoT model after $T(n)$ steps is defined as the last $m$ tokens of $f^{T(n)}(x)$, where $m$ denotes the output length.
-A.3 Looped Transformer
-Definition A.3 (Formal: Looped TF).
-Let the Transformer block be defined as the composition:
-$f\coloneq(\mathsf{id}+\mathbf{FF}_{L})\circ(\mathsf{id}+\mathbf{SA}_{L})\circ\cdots\circ(\mathsf{id}+\mathbf{FF}_{1})\circ(\mathsf{id}+\mathbf{SA}_{1}),$
-where $\mathbf{SA}_{\ell}$ and $\mathbf{FF}_{\ell}$ denote the (non-causal) self-attention and feedforward layers at depth $\ell\in[L]$.
-Given an input token sequence $x=(x_{1},\dots,x_{n})\in\mathcal{V}^{n}$, the initial hidden state is: $f^{0}(x)\coloneq\mathbf{WE}(x).$ At each loop iteration $k$, the hidden state is updated by:
-$f^{k+1}(x)\coloneq f\left(f^{k}(x)+\mathbf{PE}(\cdot,k)\right),$
-where $\mathbf{PE}(i,k)$ denotes a loop-dependent positional encoding at position $i$ and iteration $k$.
-The final output after $T(n)$ loop iterations is given by $\arg\max_{i\in[|\mathcal{V}|]}\bm{\gamma}\circ f^{T(n)}(x),$ and the model’s prediction is defined as the last $m\leq n$ tokens of this projected sequence, where $m$ denotes the output length.
-Appendix B Deferred Proofs for Section 4
-B.1 Deferred Definition
-This section provides the deferred definitions of the Assumptions 4.24.3, and 4.5.
-First, we define the notion of polynomially-efficient approximation, in which a function can be closely approximated by a two-layer ReLU network with polynomial width and logarithmic bit precision, while ensuring consistent separation between distinct outputs.
-This formulation follows the assumption in Feng et al.
-(2023), but differs in the input and output domains: whereas their setting considers real-valued functions over continuous domains, we focus on symbolic functions $f_{n}:\Sigma^{\ell(n)}\to\Sigma$ over finite alphabets.
-Accordingly, we formalize the notion as follows.
-Definition B.1 (Polynomially-Efficient Approximation).
-Let $f_{n}:\Sigma^{\ell(n)}\to\Sigma$ be a function over input strings of length $\ell(n)$, where $\Sigma=\{s_{1},\dots,s_{|\Sigma|}\}$ is a finite alphabet with fixed enumeration.
-We say that $f_{n}$ is polynomially-efficiently approximated if there exists a two-layer feedforward network with weight matrices $\mathbf{W}_{1}:\mathbb{Q}_{p(n)}^{\ell(n)\cdot|\Sigma|}\to\mathbb{Q}_{p(n)}^{m},\quad\mathbf{W}_{2}:\mathbb{Q}_{p(n)}^{m}\to\mathbb{Q}_{p(n)}^{|\Sigma|},$ where $m\in\mathrm{poly}(n)$ and $p(n)=\lceil\log_{2}n\rceil$, such that for any input $x\in\Sigma^{\ell(n)}$, the network output
-$\hat{y}(x)\coloneq\mathbf{W}_{2}\,\mathrm{ReLU}\big{(}\mathbf{W}_{1}\,\bm{\mathrm{emb}}(x)\big{)}\in\mathbb{Q}_{p(n)}^{|\Sigma|}$ (10)
-satisfies the following condition: If $f_{n}(x)=s_{j}$, then there exists a constant $\delta>0$ such that
-$\hat{y}_{j}(x)>1-\delta\quad\text{and}\quad\hat{y}_{i}(x)<1-\delta\quad\text{for all }i\neq j.$ (11)
-Here, $\bm{\mathrm{emb}}(x)\in\{0,1\}^{\ell(n)\cdot|\Sigma|}$ denotes the concatenation of the one-hot encodings of the input tokens $x=(x_{1},\ldots,x_{\ell(n)})$, where each $x_{i}=s_{k}\in\Sigma$ is represented as $\mathrm{emb}(x_{i})\coloneq\mathbf{e}_{k}\in\{0,1\}^{|\Sigma|}$.
-Then, we define a bounded fan-in computation graph as one in which each node receives input from at most a constant number of predecessors.
-Definition B.2 (Bounded Fan-In).
-The computation graph $G_{n}=(V_{n},E_{n})$ has bounded fan-in: there exists a constant $k\in\mathbb{N}$ such that for all $v\in V_{n}$, the in-degree of $v$ is at most $k$, i.e.,
-$\forall v\in V_{n},\quad\deg^{-}(v)\leq k.$ (12)
-Lastly, we define the parallel space complexity of a DAG as the maximum number of nodes that must be held in memory simultaneously during layer-wise parallel evaluation, including both the nodes in the current layer and any unresolved dependencies from earlier layers that are required to compute future layers.
-Definition B.3 (Parallel Space Complexity).
-Let $G=(V,E)$ be a directed acyclic graph. For each node $v\in V$, define its depth by:
-$\mathrm{depth}(v)\coloneq\max\left\{\ell\in\mathbb{N}\mid\exists\,u_{0}\to u_{1}\to\dots\to u_{\ell}=v\text{ in }G\right\}.$ (13)
-Then the canonical layer decomposition is given by:
-$L_{\ell}\coloneq\{v\in V\mid\mathrm{depth}(v)=\ell\}$ (14)
-for $\ell=0,1,\dots,D$, where $D\coloneq\max_{v\in V}\mathrm{depth}(v)$ is the depth of $G$.
-The parallel space complexity of $G$ is defined as
-$S(G)\coloneq\max_{\ell\in\{0,\dots,D\}}\left|L_{\ell}\cup\bigcup_{v\in L_{>\ell}}\left(\mathrm{Pred}(v)\cap\bigcup_{s\leq\ell}L_{s}\right)\right|,$ (15)
-where $L_{>\ell}\coloneq\bigcup_{s>\ell}L_{s}$, and $\mathrm{Pred}(v)\coloneq\{u\in V\mid(u\to v)\in E\}$.
-B.2 Notation and Orthogonal Vectors
-We introduce a method for constructing orthogonal vectors under finite-precision arithmetic and define the associated notation, following the results of Li et al. (2024).
-Notation
-We follow the notation from Li et al.
-(2024).
-For any positive integer $s\in\mathbb{N}^{+}$ and $x\in\{0,1,\dots,2^{s}-1\}$, we denote by $\mathsf{bin}_{s}(x)\in\{0,1\}^{s}$ the standard binary representation of $x$ using $s$ bits, defined such that $x=\sum_{i=1}^{s}2^{i}\cdot(\mathsf{bin}_{s}(x))_{i}.$ We further define the signed binary encoding of $x$, denoted by $\mathsf{sbin}_{s}(x)\in\{-1,1\}^{s}$, as $\mathsf{sbin}_{s}(x)=2\cdot\mathsf{bin}_{s}(x)-(1,\ldots,1).$ Let $x,y\in\mathbb{R}^{s}$ be two vectors of the same length.
-We define their interleaving, denoted by ${x}^{\frown}{y}\in\mathbb{R}^{2s}$, as follows: $({x}^{\frown}{y})_{2i-1}=x_{i},({x}^{\frown}{y})_{2i}=y_{i}$ for all $i\in[s].$ We denote the rounding operation by $\left[\cdot\right]_{s}\triangleq[\cdot]_{0,s}$.
-Let $\mathbf{1}_{s}$ denote the all-ones vector of length $s$.
-We also define rounding operation $\left\langle\cdot,\cdot\right\rangle_{s}$, ×$\times_{s}$, and $\sigma_{s}$ for the inner product, product, and softmax, respectively.
-To construct approximately orthogonal vectors under finite-precision inner products, Lemma E.3 of Li et al.
-(2024) shows that the vectors $q_{i}$ and $k_{j}$ are designed so that the finite-precision inner product $\langle q_{i},k_{j}\rangle_{s}$ vanishes when $i=j$, and accumulates a non-zero bias otherwise.
-This property plays a crucial role in enabling selective activation in softmax attention.
-Lemma B.4 (Li et al., 2024).
-For any $s\in\mathbb{N}^{+}$, let $\mathbf{q}_{i}={\mathsf{sbin}_{s}(i)}^{\frown}{1_{s}}$ and $\mathbf{k}_{i}=2^{s+1}\cdot({\mathsf{sbin}_{s}(i)}^{\frown}{(-1_{s})})$ for all $i\in[2^{s}-1]$, it holds that $\left[\exp(\left\langle\mathbf{q}_{i},\mathbf{k}_{j}\right\rangle_{s})\right]_{s}=\mathbf{1}[i=j]$ for all $i,j\in[2^{s}-1]$.
-B.3 Proof for Theorem 4.4
-Proof.
-Our proof follows the general strategy of Theorem 3.3 in Li et al. (2024), which analyzes CoT decoding over Boolean circuits. Here, we extend the argument to general DAG-based computations.
-Let $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$ be a directed acyclic computation graph, where each node $v\in V_{n}$ is labeled by a one-hot vector $\bm{\ell}(v)\in\{0,1\}^{|\mathcal{F}_{n}|}$ indicating the function assigned to $v$ from the finite set $\mathcal{F}_{n}$.
-Let $v_{1},\dots,v_{|V_{n}|}$ denote a fixed topological ordering of $V_{n}$, with inputs appearing first and outputs last.
-By Assumption 4.3, each node $v_{i}$ has at most $C\in\mathbb{N}$ predecessors.
-We encode the predecessor indices as a vector $\mathrm{pred}(v_{i})\in[|V_{n}|]^{C}$, with zero-padding where necessary.
-Let the vocabulary ${\mathcal{V}}=\Sigma$, and let the finite-precision parameter be $p(n)=\lceil\log_{2}n\rceil$.
-We assume that the CoT model, implemented as a single-layer decoder-only Transformer, generates the value at each node sequentially following the topological order of the computation graph.
-At the decoding step $k$, the model has access to the sequence
-$(x_{1},\,x_{2},\,\ldots,\,x_{n},\,y_{1},\,y_{2},\,\ldots,\,y_{k})\in\Sigma^{n+k},$ (16)
-where $x=(x_{1},\ldots,x_{n})\in\Sigma^{n}$ denotes the input, and $y_{i}=v_{i}(x)\in\Sigma$ denotes the value computed at node $v_{i}$ for $i\leq k$.
-We prove by induction that the model can correctly generate $y_{k+1}=v_{k+1}(x)$.
-Embedding.
-The embedding at position $i$ is a vector in $\mathbb{Q}_{p(n)}^{m}$, where $m\coloneq|{\mathcal{V}}|+|\mathcal{F}_{n}|+(1+C)p(n)+C.$ The embedding is defined as
-[…]
-where ${\bm{e}}\colon\Sigma\to\mathbb{N}^{|{\mathcal{V}}|}$ is the one-hot encoding, and $\bm{\mathrm{sbinpred}}_{p(n)}(v)\in\mathbb{Q}^{C\cdot p(n)}_{p(n)}$ encodes the binary representations of the predecessor indices:
-$\bm{\mathrm{sbinpred}}_{p(n)}(v)=\left(\mathsf{sbin}_{p(n)}(\mathrm{pred}(v)_{0}),\,\ldots,\,\mathsf{sbin}_{p(n)}(\mathrm{pred}(v)_{C-1})\right).$ (18)
-This embedding is constructed via the sum of a word embedding $\mathbf{WE}\colon{\mathcal{V}}\to\mathbb{Q}_{p(n)}^{m}$ and a positional embedding $\mathbf{PE}\colon\mathbb{N}\to\mathbb{Q}_{p(n)}^{m}$, given by:
-[…]
-[…]
-The final embedding at position $i$ is then $\mathbf{WE}(v_{i}(x))+\mathbf{PE}(i)$.
-We then provide the construction of a single-layer decoder-only Transformer.
-Attention Layer.
-The attention layer consists of $C$ heads.
-The $h$-th head is configured to attend to the position corresponding to the $h$-th predecessor of node $v_{i+1}$.
-Specifically, for each position $i\in[n+k]$ and head $h\in[C]$, the attention vectors are defined as:
-$\mathbf{q}_{i,h}={\mathsf{sbin}_{p(n)}(\mathrm{pred}(v_{i})_{h})}^{\frown}{1_{p(n)}},\,\mathbf{k}_{i,h}=2^{p(n)+1}\cdot{\mathsf{sbin}_{p(n)}(i)}^{\frown}{(-1_{p})},\,\mathbf{v}_{i,h}=({\bm{e}}(v_{i}(x)),0,\ldots,0).$ (21)
-With appropriate output projection, attention retrieves the one-hot representations of each predecessor’s value. The output at position $i$ from the attention layer is:
-$\Big{(}{\bm{e}}(v_{i}(x)),\,\bm{\ell}(v_{i+1}),\,\mathsf{sbin}_{p(n)}(i),\,\bm{\mathrm{sbinpred}}_{p(n)}(v_{i+1}),\,\underbrace{{\bm{e}}(v_{\mathrm{pred}(v_{i+1})_{1}}(x)),\,\ldots,\,{\bm{e}}(v_{\mathrm{pred}(v_{i+1})_{C}}(x))}_{\text{updated}}\Big{)}.$ (22)
-Feed-forward Layer.
-By Assumption 4.2 and Definition B.1, each function $f\in\mathcal{F}_{n}$ can be efficiently approximated by a two-layer neural network.
-Since $\mathcal{F}_{n}$ is finite, we construct a two-layer mixture-of-experts (MoE) network with one expert per function.
-For any node $v$, given the concatenated vector
-${\bm{z}}^{(v)}\coloneq\left({\bm{e}}\big{(}v_{\mathrm{pred}(v)_{1}}(x)\big{)},\,\ldots,\,{\bm{e}}\big{(}v_{\mathrm{pred}(v)_{C}}(x)\big{)}\right),$ (23)
-the feed-forward output is computed as:
-$\widehat{{\bm{e}}(v(x))}=\mathbf{W}_{2}^{(\bm{\ell}(v))}\cdot\mathrm{ReLU}\left(\mathbf{W}_{1}^{(\bm{\ell}(v))}{\bm{z}}^{(v)}\right),$ (24)
-where $\mathbf{W}_{1}^{(\bm{\ell}(v))},\mathbf{W}_{2}^{(\bm{\ell}(v))}$ are the parameters of the selected expert.
-That is, there exists a polynomial-size two-layer feed-forward network which, given input of the form form equation 22, computes an approximation $\widehat{{\bm{e}}(v_{i+1}(x))}\in\mathbb{Q}^{|\Sigma|}_{p(n)}$, where the expert is selected according to $\bm{\ell}(v_{i+1})$.
-Decoding.
-The next output symbol is then decoded as:
-$v_{i+1}(x)=s_{j}\quad\text{where}\quad j=\arg\max_{i\in[|\Sigma|]}\widehat{{\bm{e}}(v_{i+1}(x))}_{i}.$ (25)
-This selection can equivalently be expressed as an $\arg\min$ over the negated scores.
-Thus, the CoT model correctly computes $y_{k+1}=v_{k+1}(x)$, completing the inductive step.
-By induction on $k$, the model computes the values at all nodes in topological order, and the proof is complete.
-∎
-B.4 Proof for Theorem 4.7
-Generally, directed acyclic graphs (DAGs) may contain long-range dependencies that span multiple layers, complicating layer-wise simulation by Looped Transformers.
-To facilitate such simulation, it is beneficial to constrain the dependency structure so that each node depends only on nodes in the immediately preceding layer.
-To this end, we introduce a transformation called strict layered expansion, which converts any DAG into an equivalent form in which all edges connect adjacent layers by inserting identity-copy nodes that forward intermediate values.
-Definition B.5 (Strict Layered Expansion).
-Let $G=(V,E,\mathcal{F})$ be a computation graph equipped with a layer decomposition $(L_{0},\dots,L_{D})$, where each node $v\in V$ belongs to some layer $L_{\ell}$ according to its depth $\mathrm{depth}(v)=\ell$.
-The strict layered expansion of $G$, denoted by $G^{\prime}=(V^{\prime},E^{\prime},\mathcal{F}^{\prime})$, is defined as follows: For each edge $(u\to v)\in E$, where $\mathrm{depth}(u)=d_{u}$ and $\mathrm{depth}(v)=d_{v}$, insert the path
-$u^{(d_{u})}\to u^{(d_{u}+1)}\to\cdots\to u^{(d_{v}-1)}\to v^{(d_{v})}$
-with each node has identity function, into $E^{\prime}$ by introducing intermediate identity nodes as necessary.
-Then $G^{\prime}$ is strictly layered: all edges connect adjacent layers, and each dependency is limited to the immediately preceding layer.
-Importantly, this transformation preserves both the parallel space complexity (as defined in Definition B.3) and the parallel time complexity (i.e., the depth of the DAG).
-Then, we show that a Looped Transformer can simulate the strict layered expansion of a computation graph by computing each layer in a single loop iteration.
-Proof of Theorem 4.7.
-Let $G_{n}^{\prime}=(V_{n}^{\prime},E_{n}^{\prime},\mathcal{F}_{n}^{\prime})$ denote the strict layered expansion of the computation graph $G_{n}=(V_{n},E_{n},\mathcal{F}_{n})$.
-Here, $\mathcal{F}_{n}^{\prime}$ extends $\mathcal{F}_{n}$ by including the identity function, which is assigned to all newly inserted nodes.
-Define a labeling function $\ell:V_{n}^{\prime}\to\{0,1\}^{|\mathcal{F}_{n}|+1}$ that assigns to each node $v\in V_{n}^{\prime}$ the one-hot encoding of its associated function.
-By Assumption 4.5, there exists a constant $W>0$ such that the maximum layer width of $G_{n}^{\prime}$ is at most $Wn$.
-We fix a consistent ordering within each layer and pad with dummy nodes so that each layer contains exactly $Wn$ nodes.
-Let $v_{t,i}$ denote the $i$-th node in layer $t$, where $t\in\{0,\dots,D\}$ and $i\in\{1,\dots,Wn\}$, and let $v_{t,i}(x)$ denote the value it computes on input $x$.
-By Assumption 4.3, each node $v_{t,i}$ has at most $C\in\mathbb{N}$ predecessors in the previous layer $t-1$.
-We define $\mathrm{pred}(v_{t,i})\in[Wn]^{C}$ as a vector encoding the indices $j$ such that $v_{t-1,j}\in\mathrm{Pred}(v_{t,i})$, and set unused entries to 00.
-We now describe a two-layer Looped Transformer architecture that simulates the layerwise evaluation of the computation graph.
-Representation.
-Let ${\bm{h}}^{(t,i)}\in\mathbb{Q}_{p(n)}^{mW}$ be the hidden embedding at position $i\in[n]$ after the $t$-th loop iteration, where $p(n)\coloneq\lceil\log_{2}n\rceil$, and $m\coloneq|{\mathcal{V}}|+|\mathcal{F}_{n}|+1+(1+C)p(n)+C.$ Each ${\bm{h}}^{(t,i)}$ is divided into $W$ blocks of size $m$, where the $w$-th block corresponds to node $v_{t+1,\,(i-1)W+w}$, and is defined as:
-[…]
-with the components defined as:
-  • ${\bm{e}}(\cdot)\in\mathbb{N}^{|{\mathcal{V}}|}$: one-hot encoding of the computed value;
-  • $\ell(v)\in\mathbb{Q}^{|\mathcal{F}_{n}|+1}_{p(n)}$: one-hot function label;
-  • $\mathsf{sbin}_{p(n)}(j)\in\mathbb{Q}^{p(n)}_{p(n)}$: signed binary representation of index $j$;
-  • $\bm{\mathrm{sbinpred}}_{p(n)}(v)\in\mathbb{Q}^{C\cdot p(n)}_{p(n)}$: encoding of predecessor indices.
-First Layer.
-The attention layer consists of $CW$ heads indexed by pairs $(w,c)\in[W]\times[C]$.
-Each head $h=(w-1)C+c$ retrieves the $c$-th predecessor of node $v_{t+1,(i-1)W+w}$.
-The query/key/value definitions are:
-$\mathbf{q}_{i,h}$ $={\mathsf{sbin}_{p(n)}(\mathrm{pred}(v_{t+1,\,(i-1)W+w})_{c})}^{\frown}{1_{p(n)}},$ (27)
-$\mathbf{k}_{i,h}$ $=2^{p(n)+1}\cdot{\mathsf{sbin}_{p(n)}((i-1)W+w)}^{\frown}{-1_{p(n)}},$ (28)
-[…]
-After aggregation and projection, each block ${\bm{h}}^{(t,i)}_{(w)}$ is updated by writing the retrieved values into the last $C$ slots:
-$\begin{split}{\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm}\;\leftarrow\;\\ \Big{(}{\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm-C},\;&\underbrace{e(v_{\mathrm{pred}(v_{t+1,\,(i-1)W+w})_{1}}(x)),\,\ldots,\,e(v_{\mathrm{pred}(v_{t+1,\,(i-1)W+w})_{C}}(x))}_{\text{updated}}\Big{)}.\end{split}$ (30)
-Feed-forward Layer.
-By Assumption 4.2 and Definition B.1, each $f\in\mathcal{F}_{n}$ is approximable by a two-layer ReLU network.
-Since $\mathcal{F}_{n}$ is finite, a Mixture-of-Experts (MoE) architecture can be used: for each node $v$, define
-${\bm{z}}^{(v)}\coloneq\left({\bm{e}}\big{(}v_{\mathrm{pred}(v)_{1}}(x)\big{)},\,\ldots,\,{\bm{e}}\big{(}v_{\mathrm{pred}(v)_{C}}(x)\big{)}\right),$ (31)
-and compute
-$\widehat{{\bm{e}}(v(x))}={\bm{W}}_{2}^{(\ell(v))}\cdot\mathrm{ReLU}\left({\bm{W}}_{1}^{(\ell(v))}{\bm{z}}^{(v)}\right).$ (32)
-Thus, a polynomial-size two-layer FFN can compute ${\bm{e}}(v_{t+1,\,(i-1)W+w}(x))$ from equation 26. We update the embedding as:
-${\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm}\leftarrow\Big{(}\widehat{{\bm{e}}(v_{t+1,\,(i-1)W+w}(x))},\;{\bm{h}}^{(t,i)}_{1+(w-1)m+|{\mathcal{V}}|\,:\,wm}\Big{)}.$ (33)
-Second Layer.
-The attention layer is disabled (e.g., all attention weights are set to zero). To produce one-hot output, define
-$\widehat{{\bm{e}}(f_{n}(x))}_{i}\coloneq\mathrm{ReLU}\left(\hat{y}_{i}(x)-(1-\delta)\right),$ (34)
-assuming the correct index $j$ satisfies $\hat{y}_{j}(x)>1-\delta$ and all others are below that threshold. We then update:
-${\bm{h}}^{(t,i)}_{1+(w-1)m\,:\,wm}\leftarrow\Big{(}{\bm{e}}(v_{t+1,\,(i-1)W+w}(x)),\;{\bm{h}}^{(t,i)}_{1+(w-1)m+|{\mathcal{V}}|\,:\,wm}\Big{)}.$ (35)
-Positional Embedding.
-For the next iteration $t+1$, define the time-dependent positional embedding:
-[…]
-Thus, the hidden representation after the $(t+1)$-th loop matches the structure in Eq.
-equation 26.
-The output function returns the one-hot encoding of the output nodes with negated values, completing the proof.
-∎
-B.5 Proof for Theorem 4.11
-We first present the deferred definition of the class $\mathsf{SIZE}^{\mathcal{L}}$ and state the lemma showing how feedforward layers can simulate gates with unbounded fan-in.
-Definition B.6 ($\mathsf{SIZE}^{\mathcal{L}}$ (Li et al., 2024)).
-For any decision problem ${\mathcal{L}}$ and $T(n)\subseteq O(\mathsf{poly}(n))$, we define $\mathsf{SIZE}^{\mathcal{L}}(T(n))$ as the set of decision problems ${\mathcal{L}}^{\prime}$ such that there exists $p(n)\in\mathsf{poly}(n)$ and circuits $\{C_{n}\}_{n=1}^{\infty}$ where $C_{n}$ contains at most $O(T(n))$ $\mathsf{AND}$, $\mathsf{OR}$, $\mathsf{NOT}$, and ${\mathcal{L}}_{p(n)}$ gates.
-For a complexity class ${\mathcal{C}}$, we define $\mathsf{SIZE}^{{\mathcal{C}}}(T(n))\triangleq\cup_{{\mathcal{L}}\in{\mathcal{C}}}\mathsf{SIZE}^{{\mathcal{L}}}(T(n))$.
-Lemma B.7 (Li et al., 2024).
-Unlimited-fanin $\mathsf{AND},\mathsf{OR}$ (resp.
-$\mathsf{MAJORITY}):\{0,1\}^{n}\to\{0,1\}$ can be simulated by some 2-layer feedforward ReLU network with constant (resp.
-$\log{n}$) bits of precision constant hidden dimension and additional $n$ constant inputs of value 1.
-Mathematically, let $\mathbf{FF}[s(n)]$ be the set of functions $C:\{0,1\}^{n}\to\{0,1\}$ which can be a two-layer feedforward ReLU network with at most $s(n)$ bits of precision and constant hidden dimension $\mathbf{FF}_{\theta}:\{0,1\}^{2n}\to\{0,1\},\mathbf{FF}_{\theta}(x^{\prime})=\mathbf{W}_{2}\,\mathrm{ReLU}(\mathbf{W}_{1}x^{\prime}+b)$, such that for any $x\in\{0,1\}^{n}$,
-$\mathbf{FF}_{\theta}(x_{1},1,x_{2},1,\ldots,x_{n},1)=C(x).$ (37)
-We have unlimited-fanin $\mathsf{AND},\mathsf{OR}\in\mathbf{FF}[1]$ and $\mathsf{MAJORITY}\in\mathbf{FF}[\log{n}]$.
-We then provide the deferred proof for Looped Transformers.
-Proof for Theorem 4.11.
-We show that the result for $\mathsf{LOOP}[T(n),\mathrm{poly}(n),\log n]$.
-The same argument applies to the result for $\mathsf{LOOP}[T(n),\mathrm{poly}(n),1]$.
-Let $\mathsf{TC}^{[T(n)]}$ denote the class of languages decidable by a nonuniform family of threshold Boolean circuits of polynomial size and depth $T(n)$.
-$\mathsf{LOOP}[T(n),\mathsf{poly}(n),\log n]\subseteq\mathsf{TC}^{[T(n)]}$: For each loop iteration, it suffices to show that the update of each bit in each dimension of the embedding at each input position is computable by a $\mathsf{TC}^{0}$ circuit.
-This follows from the result of Merrill & Sabharwal (2023), which showed that Transformer blocks with logarithmic precision can be simulated by $\mathsf{TC}^{0}$ circuits.
-$\mathsf{TC}^{[T(n)]}\subseteq\mathsf{LOOP}[T(n),\mathsf{poly}(n),\log n]$: We show that any threshold circuit of depth $T(n)$ and polynomial size can be simulated by a $\mathsf{LOOP}[T(n),\mathsf{poly}(n),\log n]$ model.
-We construct a model in which the attention layer, in the first loop iteration, (1) aggregates the input information into a single token position.
-At each loop iteration, (2) the (looped) feedforward layer excutes the gates corresponding to each layer of the threshold circuit using Lemma B.7.
-First, we show how an attention layer combined with a feed-forward layer can be used to copy all input tokens to the final position, i.e.,$\sum_{j=1}^{n}{\bm{e}}_{j}x_{j},$ where ${\bm{e}}_{j}\in\mathbb{N}^{n}$ denotes the one-hot positional vector for index $j$, following the strategy in the proof of Theorem 3.7 in Li et al.
-(2024).
-Each input token at position $i\in[n]$ is embedded as a pair $(x_{i},{\bm{e}}_{i})$. To compute the mapping
-$(x_{i},{\bm{e}}_{i})\mapsto(x_{i},x_{i}\cdot{\bm{e}}_{i}),$ (38)
-we define a feedforward layer that takes input $(z_{0},z_{1},\dots,z_{n})$ and computes:
-$(z_{0},z_{1},\dots,z_{n})\mapsto\left(z_{0},\mathrm{ReLU}(z_{0}+z_{1}-1),\dots,\mathrm{ReLU}(z_{0}+z_{n}-1)\right).$ (39)
-To aggregate all positions via uniform attention, we use a single-head attention layer with:
-$\mathbf{q}_{i}=\mathbf{k}_{i}=\mathbf{1}_{n},\quad\mathbf{v}_{i}=nx_{i}\cdot{\bm{e}}_{i}\quad\text{for all }i\in[n],$ (40)
-where $\mathbf{1}_{n}$ is the all-ones vector in $\mathbb{R}^{n}$, and outputs become $\sum_{i=j}^{n}\frac{1}{n}\cdot\mathbf{v}_{j}=\sum_{i=j}^{n}x_{j}{\bm{e}}_{j}.$
-Then, we show that a looped feedforward layer can simulate the circuit.
-Let the threshold circuit be $G=(V,E)$, where each vertex $v_{i}\in V$ corresponds to a Boolean function $v_{i}(x)$ computed at that node for the input $x$.
-By Lemma B.7, for each node $v_{i}$, there exists a feedforward layer $\mathbf{FF}_{\theta(i)}$ s.t.
-$\mathbf{FF}_{\theta(i)}\left(v_{j_{1}}(x),1,\,v_{j_{2}}(x),1,\,\ldots,\,v_{j_{d_{i}}}(x),1\right)=v_{i}(x),$ (41)
-where $j_{1},\dots,j_{d_{i}}$ are the indices of the nodes with edges into $v_{i}$, and $d_{i}=\deg^{-}(v_{i})$.
-Then, we define a feedforward layer $\mathbf{FF}_{\theta}$ that simultaneously updates all gate values in parallel.
-Let $\ell:[|V|]\to V$ be any fixed bijection assigning an index to each gate, and define an embedding vector ${\bm{z}}\in\mathbb{R}^{2|V|}$, where each gate $\ell(i)$ is associated with a 2-dimensional slot $({\bm{z}}_{2i},{\bm{z}}_{2i+1})$.
-We use ${\bm{z}}_{2i}$ to store the current value $x_{\ell(i)}$, and fix ${\bm{z}}_{2i+1}=1$ as a constant bias input.
-The layer $\mathbf{FF}_{\theta}$ internally applies the appropriate local feedforward function $\mathbf{FF}_{\theta(\ell(i))}$ for each gate $\ell(i)$, using the current values of its input gates from the embedding.
-Formally, for each $i\in[|V|]$, the update is given by
-${\bm{z}}_{2i}\leftarrow\mathbf{FF}_{\theta(\ell(i))}\left({\bm{z}}_{2j_{1}},1,{\bm{z}}_{2j_{2}},1,\ldots,{\bm{z}}_{2j_{d_{i}}},1\right),$ (42)
-where $\ell(j_{1}),\dots,\ell(j_{d_{i}})$ are the gates with edges into $\ell(i)$.
-After $T(n)$ loop iterations—where $T(n)$ is the depth of the circuit DAG—the embedding stores the correct output value $x_{v}$ for every gate $v\in V$.
-By appropriately combining the components described above, we construct a $3$-layer Transformer that computes the desired operation.
-We set the embedding size to be $m=1+n+2|V|$.
-Given an input sequence $(x_{1},x_{2},\dots,x_{n})\in\{0,1\}^{n}$, each input token $i\in[n]$ is mapped, in the token and positional embedding step, to the initial representation
-[…]
-where ${\bm{e}}_{i}\in\{0,1\}^{n}$ denotes the one-hot encoding of position $i$, and $\mathbf{0}_{2|V|}\in\mathbb{R}^{2|V|}$ is a zero vector.
-For each loop iteration $k$, given the input from the previous iteration in the form
-$(x_{i},\,{\bm{e}}_{i},\,h_{1}^{(k)},\,h_{2}^{(k)},\,\dots,\,h_{n+2|V|}^{(k)}),$ (44)
-the model processes the input and applies computations across layers as follows: In the first two layers, the model extracts and aggregates the input values $x_{1},\dots,x_{n}$.
-Then, in the third layer, the feedforward layer computes the circuits.
-We describe the computation for each layer as follows:
-  1. 1.
-    First layer: The attention layer is disabled (e.g., all attention weights are set to zero), so the input is passed through unchanged.
-The feed-forward layer computes $x_{i}{\bm{e}}_{i}$ for each position:
-    $(x_{i},\,\underbrace{x_{i}{\bm{e}}_{i}}_{\text{updated}},\,\mathbf{0}_{2}n,\,h_{2n+1}^{(k)},\,\dots,\,h_{n+2|V|}^{(k)}),$ (45)
-  2. 2.
-    Second layer: The attention layer aggregates the $x_{i}{\bm{e}}_{i}$ values into the last token position using uniform attention, as described earlier, producing:
-    $(x_{i},\,x_{i}{\bm{e}}_{i},\,\underbrace{x_{1},\,1,\,\dots,\,x_{n},\,1}_{\text{updated}},\,h_{2n+1}^{(k)},\,\dots,\,h_{n+2|V|}^{(k)}).$ (46)
-    This ensures that the input vector $x=(x_{1},\dots,x_{n})$ is correctly stored and made available at every subsequent iteration. (The feed-forward layer is disabled.)
-  3. 3.
-    Third layer: The attention layer is disabled, and the feed-forward layer begins simulating the desired gate computations as defined in Eq.
-42, where the input nodes correspond to the first $n$ components of the embedding.
-Finally, the output function selects the value of the output gate. This completes the construction. ∎
-B.6 Definitions and Proofs for Proposition 4.14
-We introduce deferred definitions for each problem and provide proofs showing that each problem lies in nonuniform $\mathsf{NC}^{2}$, thereby proving the statement.
-Definition B.8 (Solving Integer Linear Systems).
-Given an $n\times n$ integer matrix $A\in\mathbb{Z}^{n\times n}$ and a vector $b\in\mathbb{Z}^{n}$, the task is to determine whether there exists a solution $x\in\mathbb{Q}^{n}$ to the system of linear equations $Ax=b$, and to compute such a solution if it exists.
-Proposition B.9.
-The problem of solving integer linear systems over $\mathbb{Q}$, where $A\in\mathbb{Z}^{n\times n}$ is non-singular, lies in $\mathsf{NC}^{2}$.
-Proof Sketch.
-Using Csanky’s algorithm, the inverse of an $n\times n$ matrix over the rationals (or over a field of characteristic zero) can be computed in $\mathsf{NC}^{2}$ (Csanky, 1975).
-Once $A^{-1}$ is computed, the solution is given by $x=A^{-1}b$, which is a matrix-vector multiplication and hence also lies in $\mathsf{NC}^{1}$.
-Therefore, the full solution procedure belongs to $\mathsf{NC}^{2}$.
-∎
-Definition B.10 (Perfect Matching in Bipartite Graphs).
-Given a bipartite graph $G=(U\cup V,E)$, where $|U|=|V|=n$, decide whether $G$ contains a perfect matching—that is, a subset of edges $M\subseteq E$ such that every vertex in $U$ and $V$ is incident to exactly one edge in $M$.
-Proposition B.11.
-The bipartite perfect matching problem is in nonuniform $\mathsf{NC}^{2}$.
-Proof Sketch.
-The problem can be reduced to checking whether a certain symbolic matrix has a non-zero determinant, and the determinant of an $n\times n$ matrix can be computed in randomized uniform $\mathsf{NC}^{2}$ (Karp et al., 1985).
-By Adleman’s theorem (Adleman, 1978), any $\mathsf{BPP}$ algorithm can be simulated in $\mathsf{P}/\mathsf{poly}$, meaning that for each input length $n$, the randomness used by the algorithm can be fixed in advance and hard-wired as advice.
-This yields a nonuniform circuit family of depth $O(\log^{2}n)$ and polynomial size.
-Thus, the problem lies in nonuniform $\mathsf{NC}^{2}$.
-∎
-Definition B.12 (Fixed Context-Free Language Recognition).
-Let $G$ be a fixed context-free grammar. Given an input string $w\in\Sigma^{*}$, decide whether $w\in L(G)$, i.e., whether $w$ is generated by $G$.
-Note that this is the nonuniform word problem, and is distinct from the uniform CFG membership problem, where the grammar is part of the input and the problem is known to be $\mathsf{P}$-complete.
-Proposition B.13 (Ruzzo, 1979).
-The word problem for a fixed context-free language is in $\mathsf{NC}^{2}$.
-Appendix C Background on Approximate Counting and Sampling
-This section provides an overview of key definitions, notation, and foundational results concerning approximate counting and almost uniform generation, with a particular focus on self-reducible relations.
-C.1 Definitions of Approximate Counting
-Many computational problems can be formalized as counting the number of solutions satisfying a certain predicate.
-Formally, a counting problem is defined for relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$, where $\Sigma$ is a finite alphabet.
-For each input $x\in\Sigma^{*}$, the goal is to compute the number of feasible solutions $y\in\Sigma^{*}$ such that $(x,y)\in R$; that is, counting function #$\#R:\Sigma^{*}\to\mathbb{N}$ is defined by #$\#R(x)\coloneq\left|\left\{y\in\Sigma^{*}\mid(x,y)\in R\right\}\right|.$
-The complexity class #$\#\mathsf{P}$, introduced by Valiant (1979), captures the class of functions that count the number of accepting paths of a nondeterministic polynomial-time Turing machine.
-Definition C.1 (#$\#\mathsf{P}$).
-A function $f:\Sigma^{*}\to\mathbb{N}$ belongs to the class #$\#\mathsf{P}$ if there exists a polynomial-time Turing machine $M$ and a polynomial $p$ such that, for every input $x\in\Sigma^{*}$,
-$f(x)=\left|\left\{y\in\{0,1\}^{p(|x|)}\mid M(x,y)=1\right\}\right|.$ (47)
-Intuitively, #$\#\mathsf{P}$ encompasses problems such as counting the number of satisfying assignments of a Boolean formula (#$\#\mathrm{SAT}$), computing the permanent of a 0,1 matrix, or counting the number of perfect matchings in a bipartite graph (Valiant, 1979).
-All of these are known to be #$\#\mathsf{P}$-complete.
-Approximate Counting
-Since exact computation of #$\#\mathsf{P}$ functions is believed to be intractable, much attention has been devoted to the design of efficient approximation algorithms.
-This motivates the study of fully polynomial randomized approximation schemes (FPRAS) and their connections to approximate uniform generation, which we describe in the subsequent sections.
-Definition C.2 (FPTAS).
-Let $R$ be a relation, and define $f(x)\coloneq\lvert R(x)\rvert$.
-A fully polynomial-time approximation scheme (FPTAS) is a deterministic algorithm $A(x,\varepsilon)$ that, for any input $x\in\{0,1\}^{*}$ and any error tolerance $\varepsilon>0$, produces a value $\hat{f}(x)\in\mathbb{Q}$ such that
-$(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$ (48)
-and the running time of $A$ is bounded by a polynomial in both the input size $|x|$ and $1/\varepsilon$; i.e., $\mathrm{poly}(|x|,1/\varepsilon)$.
-Let the class $\mathsf{FPTAS}$ denote the set of all functions that admit such a scheme.
-Randomized Schemes for Approximation
-While an FPTAS provides a deterministic polynomial-time algorithm for approximating counting functions within any desired relative error, such schemes are known to exist only for a very restricted class of problems.
-The concept of randomized approximation schemes was introduced by Karp & Luby (1983), who proposed using Monte Carlo algorithms to approximate the value of hard counting functions.
-Their work laid the foundation for what is now known as the fully polynomial randomized approximation scheme (FPRAS), which allows one to efficiently obtain high-probability multiplicative approximations for a broad class of #$\#\mathsf{P}$ functions.
-Definition C.3 (FPRAS).
-Let $R$ be a relation, and define $f(x)\coloneq\lvert R(x)\rvert$.
-A randomized approximation scheme is an algorithm that, given any error tolerance $\varepsilon>0$ and failure probability $\delta>0$, outputs an estimate $\hat{f}(x)$ such that
-$\Pr\left[(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x)\right]\geq 1-\delta.$ (49)
-If the algorithm runs in time polynomial in $|x|$, $1/\varepsilon$, and $\log(1/\delta)$, it is called a Fully Polynomial-Time Randomized Approximation Scheme (FPRAS) for $f$.
-We denote by $\mathsf{FPRAS}$ the class of functions that admit such an algorithm.
-C.2 Complexity of Self-Reducible Problems
-The concept of self-reducibility, introduced by Schnorr (1976), plays a central role in bridging approximate counting and approximate sampling.
-Informally, a relation $R(x)$ is self-reducible if any solution $y\in R(x)$ can be obtained by recursively solving a polynomial number of smaller subproblems derived from $x$.
-Definition C.4 (Self-Reducibility (Schnorr, 1976)).
-A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is self-reducible if:
-  1. 1.
-    There exists a polynomial-time computable function $g\in\Sigma^{*}\to\mathbb{N}$ such that $xRy\Rightarrow|y|=g(x);$
-  2. 2.
-    There exist polynomial-time computable functions $\psi\in\Sigma^{*}\times\Sigma^{*}\to\Sigma^{*}$ and $\sigma\in\Sigma^{*}\to\mathbb{N}$ s.t.
-    $\sigma(x)$ $=O(\log|x|),$ (50)
-    $g(x)>0$ $\Rightarrow\sigma(x)>0\quad\forall x\in\Sigma^{*},$ (51)
-    $|\psi(x,w)|$ $\leq|x|\quad\forall x,w\in\Sigma^{*},$ (52)
-    and such that, for all $x\in\Sigma^{*}$, $y=y_{1}\dots y_{n}\in\Sigma^{*}$,
-    $\langle x,y_{1},\dots,y_{n}\rangle\in R\iff\langle\psi(x,y_{1}\dots y_{\sigma(x)}),y_{\sigma(x)+1},\dots,y_{n}\rangle\in R.$ (53)
-Self-reducibility relation bridges the gap between approximate counting and almost uniform generation.
-A fundamental result by Jerrum et al.
-(1986) shows that, under the self-reducibility condition, these two problems have the same time complexity: the ability to approximately count solutions within a polynomial factor implies the ability to sample from an almost uniform distribution over the solution set, and vice versa.
-Theorem C.5 (Jerrum et al., 1986).
-Let $R\subseteq\Sigma^{*}\times\Sigma^{*}$ be a self-reducible relation. There exists an FPRAS for approximating $|R(x)|$ if and only if there exists an FPAUS for sampling uniformly from $R(x)$.
-Appendix D Deferred Proofs for Section 5
-D.1 Proof for Theorem 5.6
-Proof.
-By Lemma 5.5, a probabilistic CoT model can simulate a polynomial-time probabilistic Turing machine, and hence can implement any FPRAS algorithm within polynomial CoT steps.
-In contrast, deterministic Looped Transformers with a polynomial number of loop iterations correspond to deterministic polynomial-time Turing machines, and hence can only compute functions in $\mathsf{FPTAS}$, assuming that each loop iteration performs a fixed computation deterministically.
-Given the assumption that $\mathsf{FPTAS}\neq\mathsf{FPRAS}$, there exists a relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ such that approximating $|R(x)|$ to within a relative error $\varepsilon>0$ lies in $\mathsf{FPRAS}\setminus\mathsf{FPTAS}$.
-Therefore, a probabilistic CoT model can approximate $|R(x)|$ within arbitrary relative error with high probability in polynomial time, whereas no deterministic Looped Transformer with polynomial loop iterations can achieve the same approximation.
-∎
-D.2 Proof for Lemma 5.11
-Proof.
-Fix any token position $i\in[m]$. By the definition of $(\alpha,\gamma)$-weak token-wise learning, we have
-$\Pr_{r_{i}\sim\pi(r_{i}\mid x,y_{<i})}\left[\mathcal{L}_{\mathrm{CE}}(x,y_{<i},r_{i})\leq\alpha(|x|)\right]\geq\tfrac{1}{2}+\gamma.$ (54)
-Since $\mathrm{KL}(p\,\|\,\pi)\leq\mathcal{L}_{\mathrm{CE}},$ it follows that
-$\Pr_{r_{i}}\left[\mathrm{KL}\left(p(\cdot\mid x,y_{<i})\,\|\,\pi(\cdot\mid x,y_{<i},r_{i})\right)\leq\alpha(|x|)\right]\geq\tfrac{1}{2}+\gamma.$ (55)
-By Pinsker’s inequality, this implies
-$\Pr_{r_{i}}\left[\|\pi^{(t)}-p\|_{\mathrm{TV}}\leq\sqrt{\tfrac{1}{2}\alpha(|x|)}\right]\geq\tfrac{1}{2}+\gamma.$ (56)
-From the definition of total variation distance,
-$\|\pi^{(t)}-p\|_{\mathrm{TV}}=\max_{A\subseteq\mathcal{V}}\left|\pi^{(t)}(A)-p(A)\right|\geq\left|\pi^{(t)}(y)-p(y)\right|\quad\text{for all }y\in\mathcal{V}.$
-Hence,
-$\Pr_{r_{i}}\left[\left|\pi^{(t)}(y)-p(y)\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\right]\geq\tfrac{1}{2}+\gamma.$ (57)
-Let $Z_{t}=\mathbf{1}\left[\left|\pi^{(t)}(y)-p(y)\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\right]$, and note that $\mathbb{E}[Z_{t}]\geq\tfrac{1}{2}+\gamma$.
-Let $S=\sum_{t=1}^{T}Z_{t}$ denote the number of good samples.
-Then the expectation satisfies $\mathbb{E}[S]=\sum_{t=1}^{T}\mathbb{E}[Z_{t}]\geq T\left(\tfrac{1}{2}+\gamma\right).$
-By a standard Chernoff bound:
-$\Pr\left[S<(1-\varepsilon)\mathbb{E}[S]\right]\leq\exp\left(-\frac{\varepsilon^{2}\mathbb{E}[S]}{2}\right).$ (58)
-To bound the probability that fewer than half the samples are good, we set
-$(1-\varepsilon)\mathbb{E}[S]=\tfrac{T}{2}\quad\Rightarrow\quad\varepsilon=1-\frac{T/2}{\mathbb{E}[S]}\geq\frac{2\gamma}{1+2\gamma}\geq\gamma,$ (59)
-where the last inequality holds for $\gamma<1/2$.
-Substituting into the Chernoff bound gives
-$\Pr\left[\sum_{t=1}^{T}Z_{t}<\tfrac{T}{2}\right]\leq\exp\left(-\frac{\gamma^{2}\cdot\mathbb{E}[S]}{2}\right)\leq\exp(-2T\gamma^{2}),$ (60)
-where the final inequality uses $\mathbb{E}[S]\geq T(\tfrac{1}{2}+\gamma)$.
-Therefore, if $T\geq\frac{2}{\gamma^{2}}\log\left(\frac{|\mathcal{V}|}{\delta}\right)$, then with probability at least $1-\delta/|\mathcal{V}|$, more than half of the values $\pi^{(t)}(y)$ lie within $O\left(\sqrt{\alpha(|x|)}\right)$ of $p(y)$.
-Thus, the median $\pi^{\star}(y)$ also lies within this range.
-Applying a union bound over all $y\in\mathcal{V}$, we conclude that with probability at least $1-\delta$,
-$\left|\pi^{\star}(y)-p(y)\right|\leq O\left(\sqrt{\alpha(|x|)}\right)\quad\text{for all }y\in\mathcal{V}.$ (61)
-D.3 Proof for Theorem 5.13
-Proof.
-The proof follows the approach of rejection sampling based on approximate next-token predictions, inspired by Jerrum et al. (1986).
-Let $R(x)\coloneqq\{y\in\Sigma^{*}\mid(x,y)\in R\}$ denote the solution set for input $x\in\Sigma^{*}$, and let $p(\cdot\mid x)$ be the uniform distribution over $R(x)$.
-By the self-reducibility of $R$, there exists a polynomially bounded function $m:\mathbb{N}\to\mathbb{N}$ such that all $y\in R(x)$ satisfy $|y|\leq m(|x|)$.
-Assume the CoT model $\pi$ satisfies $(\alpha,\gamma)$-weak token-wise learning with respect to $p$, for some constant $\gamma\in(0,\tfrac{1}{2})$ and function $\alpha:\mathbb{N}\to\mathbb{R}_{>0}$.
-Then for each token position $i\in[m(|x|)]$, with probability at least $1/2+\gamma$, the prediction $\pi(y_{i}\mid x,y_{<i})$ approximates the true conditional probability $p(y_{i}\mid x,y_{<i})$ within total variation distance $\leq 1/\alpha(|x|)$.
-By Pinsker’s inequality, this implies the pointwise deviation:
-$\left|\pi(y_{i}\mid x,y_{<i})-p(y_{i}\mid x,y_{<i})\right|\leq\sqrt{\frac{1}{2\alpha(|x|)}}.$ (62)
-Since $p(y_{i}\mid x,y_{<i})\geq 1/|R(x)|$, we have the relative error bound:
-$\left|\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}-1\right|\leq\sqrt{\frac{1}{2\alpha(|x|)}}\cdot|R(x)|.$ (63)
-Choose $\alpha(|x|)\geq\frac{1}{4\cdot m(|x|)^{2}\cdot|R(x)|^{2}}$, so that:
-$\left|\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}-1\right|\leq\frac{1}{2m(|x|)}.$ (64)
-Then for all $y\in R(x)$, we have:
-$\frac{\pi(y\mid x)}{p(y\mid x)}=\prod_{i=1}^{m(|x|)}\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}\in\left[\left(1-\frac{1}{2m(|x|)}\right)^{m(|x|)},\left(1+\frac{1}{2m(|x|)}\right)^{m(|x|)}\right].$ (65)
-Using the inequality $\left(1+\tfrac{1}{2m}\right)^{m}\leq e^{1/2}$, we obtain:
-$\frac{1}{\beta}\leq\frac{\pi(y\mid x)}{p(y\mid x)}\leq\beta,\quad\text{where }\beta=e^{1/2}.$ (66)
-Now define the rejection sampling algorithm:
-  • Sample $y\sim\pi(\cdot\mid x)$;
-  • If $y\notin R(x)$, reject and repeat;
-  • If $y\in R(x)$, accept with probability $\frac{1}{\beta}\cdot\frac{p(y\mid x)}{\pi(y\mid x)}\leq 1$;
-This produces exact samples from $p(\cdot\mid x)$. For any $y\in R(x)$, the probability of acceptance is:
-$\pi(y\mid x)\cdot\left(\frac{1}{\beta}\cdot\frac{p(y\mid x)}{\pi(y\mid x)}\right)=\frac{1}{\beta|R(x)|},$ (67)
-which is independent of $y$. Hence, the output is uniform over $R(x)$.
-Error Analysis.
-There are two sources of error:
-  1. 1.
-    The condition in Eq. 62 may fail to hold at some positions $i\in[m(|x|)]$.
-  2. 2.
-    The rejection sampling procedure may fail to produce any accepted sample within a bounded number of trials.
-To address (1), we use the Lemma 5.11 and draw $T=O\left(\frac{1}{\gamma^{2}}\log\left(\frac{|\mathcal{V}|\cdot m(|x|)}{\varepsilon/2}\right)\right)$ independent reasoning paths to estimate each conditional probability.
-By a union bound over all positions and tokens, the probability that any of them exceeds the deviation bound is at most $\varepsilon/2$.
-To address (2), note that the acceptance probability is at least $1/\beta^{2}$.
-Hence, by running $N=O\left(\beta^{2}\log\left(\frac{2}{\varepsilon}\right)\right)$ independent trials, the probability that all of them are rejected is at most
-$(1-1/\beta^{2})^{N}\leq e^{-N/\beta^{2}}\leq\varepsilon/2.$ (68)
-Combining both sources of error, with probability at least $1-\varepsilon$, the algorithm outputs a sample from a distribution $q^{\pi}_{\varepsilon}(\cdot\mid x)$ satisfying
-$\bigl{\|}q^{\pi}_{\varepsilon}(\cdot\mid x)-p(\cdot\mid x)\bigr{\|}_{\mathrm{TV}}\leq\varepsilon.$ (69)
-Running Time.
-Each token prediction and rejection trial takes polynomial time in $|x|$, and the number of trials is bounded by $O(\log(1/\varepsilon))$ with high probability.
-Therefore, the total running time is polynomial in $(|x|,\log(1/\varepsilon))$.
-∎